A Complete Bibliography of Publications in Applied Mathematics and Computation: 1990–1994

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

20 December 2023 Version 1.06

Title word cross-reference

(r, s) [PSMM93]. 3 [Hei91, Jin94]. ^{2} [IST93]. AX + XB = C [SH92]. c [Mit92]. D [Pat92]. e^{nx} [Kin91]. F [MP90]. GV ψ_{0} [Kin94]. h [IST93]. H^{+} [BMS92]. K^{+} [BMS92]. L [GG91b]. L^{2} [Rap92]. L^{p} [GK94]. L_{1} [dK94]. L_{p} [Sen93]. M^{[x]}/C_{2}/S/N [MES94]. N [Cha93, SMO91, SV92b]. n ≥ 0 [Kin91]. O(n^{2}) [Wai90]. P [Sim92, FLG92]. Q [CDW91, LS93b]. S [Sav93]. x ≤ 0 [Kin91]. Y′′ = f(t, Y) [JMV93].

-analysis [CDW91]. -approximation [dK94]. -bit [Cha93]. -cycle [MP90]. -D [Jin94]. -directed [PSMM93]. -expansion [IST93]. -Markov [LS93b]. -node [Pat92]. -person [SMO91]. -solvability [GK94].

-stable [Sim92]. -system [Sav93]. -systems [GG91b]. -type [MS94c]. -version [Mit92].

1994 [Cab95].

2 [Ada93b]. 2-D [CXL94, LYX^{+}93].

63 [Cab95]. 65th [Kag91a].

abstract [Opp92]. abstraction [SHR94]. Acceleration [Bar91, BCG91, Cha90a]. acceleration-pressure [Cha90a]. accelerator [CK90]. accidentally [Ten91]. accuracy [Cic93b, Sar93]. accurate [BTW94, MEZI93, PR93]. across [BMS92]. active [YF91]. actuators [Bal94]. ADA [Hus90]. Adaptive [IBS94, AH94, And90, FLYD94, KT91, Swo91, SY93, YS94, Zas93]. additive [MDR94]. adjoint [MSLB92]. aeromedical [MFS93]. aesthetics [Mur94]. aggregation [AB94]. AI [Bad93]. aided [TCS92]. AIDS [Yak92b]. air [Sla93a]. airfoil [Gol93]. Algebraic [NG94, Bai91, Ben92, CWL92, Kha91b, Koč92, Sav93, Sen90, TGK91, Zas93, Hua91]. algorithm [AH94, AR93, Bar91, BG93, EA94, FLYD94, KMR^{+}91, LAK94, LSS94, LSVS94, LW90, Mor90, Mor94, Pes91, Sch92, SW94, SHR94, Wai90, WJ92, Yak92a, Yan94, ZW90a]. algorithmic [AS90, AS92]. Algorithms [DHK90, ADM91, Cha91, GT91, HL91, Kwo93, SH92, Szi90, Sch91b]. alkyl [KK91]. all-time-step [RTJL92]. allocation [Dro90]. almost [Sim92]. ALOHA [Sch92]. along [CFT^{+}93]. Alternating [Jin94, Ela94]. Alternative [EB91, GT91, Ji94]. alternatives [MDR94]. among [LL94]. amplitudes [Lan92]. Analysis [HB91, JV90, MGL92, NW90, ALMS90, BBD91, CW92, CXL94, CDW91, CK90, CJF93, Con93c, EA94, El 92, Eno90, FL94, GP94, HJ92, HF91, IBS94, Kon92, Kum92b, KR91, LS91a, MFS93, MU90a, MU90b, Mei93, Pap94, SS92, SPD93, SHR94, Tes91b, VWSS94, Wu94, YF91]. Analytic [JV94, MN91, Ioa92, ZW90a]. Analyzing [HKLS93]. Answering [HKLS93]. any [Ada93b]. Appel [MS93b]. Application [SS92, Yee93, Cai93, Ji94, KN90, KK90, Kum92b, Mur91b, PY94, SP94, Unn90]. Applications [Cic92, Eva90a, ER94b, ER94a, LS91a, BH93, DGP93, SYS94]. Applied [Cab95, Faz90, Kag91b, Mus94, Ano92a]. approach [AaHM93, BD94, BCG91, Ezz94, FLZ92, LSP91, Mad90, Rap92, RTJL92, RR93, Uen91, Yip92]. Approximate [Kin93, She90, Jód92, KCS90, SR90, SS94a]. approximating [Ada93b]. approximation [DDL93, Fun91, Jac90, Jol92, KN90, Kum92b, MS93a, MS93b, Opp92, Sar93, Yan94, dK94]. approximations [Ham94a, Jac92, Pat91, RWB94]. arbitrary [GS94b, Wan94]. arc [CK90]. architectures [El 92]. arguments [LYC93]. arising [PA94]. arithmetic [ZW90b]. ARMA [CC94]. armature [CK90]. arrays [DGV94]. arrivals [MES92]. Artificial [GKn91, BTW94, KG91]. aspects [Bog91, Ham92, MO92]. assessed [MDR94]. Assessment [Bha94]. Assignment [MTP93]. assisted [Kor92]. associative [KKU91, KKM90]. assumptions [Arg94, CWL92]. Asymptotic [GS94b, GL94, Liq91, PSMM93, Rua92, VB92, Zag92, CC94]. asymptotical [LS91b]. asymptotics [Kha90, Kha91a, Sch92]. Atkinson [GC94]. atmosphere [Ten90]. attractivity [PPS94]. attribute [MDR94, YLS93]. attributes [YLS93]. autocorrelations [CC94]. automata [Ada93a, KK90, Kub94a, Kub94b]. automated [Tes91b]. Automatic [SW92, Hus90]. automation [CDW91]. automaton [KK91, NG94, YGH90]. Autonomic [Gun94]. autonomous [IST93, Jia94, KG91]. autonomously [GKn91]. autonomy [Gun93, Shi93]. available [MTP93]. average [Sch92]. Averages [HKLS93]. Averaging [HKLS93]. avoiding [JL93]. away [BD90]. axisymmetric [Ade94, HF93].

B [Ham94a]. B-spline [Ham94a]. balance [Den90]. ball [ZW90a]. Banach [Arg90, AC94, CAQ93, Liu94]. bandit [PY94]. bandlimited [Kha92]. bar [PZ94]. based [Bad93, Bar91, GSK93, KM94, Kon92, Mos90, PL93, Sla91]. basis [IPP91, RK93, ZM91]. Bayesian [AB94]. beam [Lan92]. behavior [ESS94, ES94, Ela94, ES90, GS94b, Zag92]. Being [Mat93]. Bellman [ZED91]. Bessel [RK93, Sri94]. Best [Kag91a, TM92]. Beyond [Mur92]. Bezout [Wam92]. bias [Jac93]. biasing [Cic93a]. biattribute [SPD93]. BIBO [ZXY94]. Bicriterion [PL93]. Bifurcation [Cas94, PBL92, ASZ94, KS92a, KR91, Nam90a, Nam90b]. binary [DGV94]. Bingham [Han92]. binomial [Jol92]. binpacking [Dro90]. Biochemical [TW90]. Bioeconomic [Gar90]. Biological [Shi93]. biology [Mat93]. biomass [BD90, Ten90]. biomass-carbon [Ten90]. biomolecule [Ten91]. Biophysicalism [Con93b]. birthday [Kag91a]. bit [Cha93]. block [Jin94]. board [Ano90b, Ano90e, Ano91a, Ano91b, Ano91d, Ano92b, Ano93b, Ano93d, Ano94e, Ano94g, Ano90a, Ano90c, Ano90d, Ano90f, Ano91c, Ano92c, Ano92d, Ano92e, Ano93a, Ano93c, Ano94a, Ano94b, Ano94c, Ano94d, Ano94f]. Bob [Cas91, Kag91a]. bodies [GGR91, WN91]. Boolean [Mit92, MO93]. Born [Kin94]. boundaries [GKn91]. Boundary [EL90, AS94a, AS94b, AR93, Bai91, CN90, Cab94, Cab95, Cas94, CM90, Faz90, GM94, GC94, Gup90, GK94, HM94, HSZ94, IPP91, JNM92, KB93, KS93b, LMS92, LM94, ML90, MFF92, OT92, OT93, PA94, Raj94, SV94, ŠNG92, SS94a, Ste94, TM92, WJ92]. boundary-layer [AS94a, HM94]. boundary-value [Bai91, Faz90, Gup90, GK94, KB93, ML90, SS94a]. Bounded [Liu94]. bounding [ZW90a]. bounds [Jód92]. Boussinesq [MEZI93, Siv92]. Boussinesq-type [MEZI93]. Brinkman [BH93, HB93, HB91]. buoyancy [KS93a]. buoyancy-induced [KS93a]. Burgers [RC91]. BVPs [Lop92].

Cahn [Fre90]. Calculation [Lan92, BTW94, JM93, VCM91]. calculations [Wam92, Zas93]. Calculus [TGK91, MU90a, MU90b]. Cantor [Bad93]. capacitated [She90]. capital [Chu92]. capture [Yip93]. carbon [Ten90]. Carlo [RBR94]. cascades [JM93]. cases [War94]. Cauchy [Sen93, Uen91]. CDF [Ahm93]. cell [And90, BMS92]. cells [CFT^{+}93]. Cellular [Kub94b, YGH90, Ada93a, KK90, KK91, Kub94a, NG94]. cellular-automaton [KK91]. cerebellar [CFT^{+}93]. certain [AS94a, CS93]. certainty [Mal93]. ceteris [Mat93]. chain [KK91, LF94, Wai90, Zag91]. channel [Ham94b]. channels [Ham94b]. Chaos [Ebe93, TW90]. Chaotic [Wer93, FG94b]. Characterization [FEGL90]. characterized [SW91]. Chebyshev [KS93a, PS94a]. chemical [Pat90]. Chen [Arg94]. chiral [TW90, Ten91]. cholesterol [Kub94b]. circuit [PS94a]. circular [enNAa94, HF93]. circumscribed [NP92]. Class [Red90b, GS94a, GM93, HSZ94, JN91, KS93b, Pat92, Phi92, TFM93]. classic [HS94]. classical [FC93, KS93b, RS92, RBR94]. classification [Wu94]. Closed [NFJ94, JNM92]. closed-form [JNM92]. cluster [MGL92]. cochlear [YF91]. code [Bar91]. Coefficient [MS92b, Bha94, IPP91]. Coefficient-parameter [MS92b]. coefficients [Ela94, Eno90, MMP90, RBA92, SS94a]. cognition [Pat93]. coherence [FE90a]. coincidence [FE90b]. colleague [Cas91]. collision [Sch92]. collocated [Bal94]. collocation [Bai91, GL94, KS93a, Liq91]. combination [Hoj94, ZP92]. combustion [Bar91]. compact [BG93]. Comparison [AS90, Cic93a, RBR94, HV90, Sch91a, VS90]. compartmental [Pat91]. competition [Zag92]. competitive [LSP91, SP93]. complementary [Moh90, Zag92]. complete [RK93]. complex [Eva90a, EHE93, ER94b, ER94a]. Complexity [NP92, Cha92c, Cha92d, Cha92a, Cha92e, Cha93, FS94]. composed [Kin91]. compressibility [BTW94]. compressible [VWSS94]. Comput [Cab95]. Computation [Ioa92, Con93c, GG91b, HJ92, How94, KS93a, Kha91a, Kha94, LW90, MO93, Sem92, Ano92a]. Computational [BS93, CW91, CW92, Ham92, ZG90, Bad93, Cic93a, HLW94, JR93, Sel94, SS94b, dK94]. Computationally [Gro93]. Computations [Fre90]. Computer [Raj94, Kor92, PZ94, TCS92, Zhe93]. computer-aided [TCS92]. computer-assisted [Kor92]. computers [Rug93, ZP92, Zhu94]. Computing [CC94, MS92c, FLZ92]. concentrating [WM91]. concept [Kai91, Sch91a]. Concise [LSS94]. condition [CM90, Gup90]. conditioning [GB90, ML90]. conditions [AR93, Cab94, Cab95, FLZ91, Gyö90, Ham94b, Lin94, Ram90, SP93, WJ92]. conductivity [Kub94b]. cones [Gyö90]. confined [CJF93]. conflict [HF91]. Conjecture [HKLS93]. conjugate [AH94, FT94, HV90]. connected [WN91]. connecting [DFM94]. conservation [Jim94, Li94]. conserving [FV91]. constant [SW91]. constitutes [Gun93]. constrained [CUK92, Flå92, Ge90a, KU92, RTJL92, TLG90, Yak92a]. constraints [Shi93, Vev92]. Construction [Ham94a, KG91]. constructive [Con93c]. consumer [SM94]. consumption [Pap92]. contact [Dav94]. continuation [KT91, MM91, MS92c, MS92b, O’N90, VBH91]. continuous [Pap94, Pat91]. continuous-time [Pat91]. contractions [LS91b]. Control [And93, And90, Avg94, BD90, Bal94, Bog91, Chu94, NA91, Pat93, Pic92, RTJL92, Sag91, Son93, Spi90, Swo91, SP93, TLG90, Udw91, Udw92a]. Controllability [Sos91, CUK92, Chu92, SY93]. controller [TFM93]. convection [Ezz94, HM94, Yee93]. convection-radiation [HM94]. Convergence [EA94, HKLS93, Hua91, UK94, AS92, Arg94, BCG91, Cro93, HJ92, Szi90, Udw92b]. convergent [BJQBQ94, SV92a]. converging [BG93]. convex [Sor93, Sor94, Sor97]. convexized [GQ90]. convolution [Jol92, RRA90]. cooling [Abd92]. Cooperation [HF91]. core [Zas93]. correction [Bar91, JNF95]. corrector [Kha91a]. correlation [Bha94]. Cost [Dro90, MW93, Pic92]. Couette [Uen91]. counting [FE90b]. coupled [JL92, JN91, JF91, JV94, Mei94]. covariance [CC94, LS93b]. cover [RBR94]. Coxian [MES92]. Crank [Jin94]. created [Ten91]. creation [Shi93]. CRI [Sch92]. Criteria [FLG92, ES93b, Mal93]. crossing [PY94]. Cubic [Ste94, Beh92, BG93, KB93, Ram90, Wan94]. current [CXL94, MS94a]. curved [Ham92]. curves [ASZ94, KS92a, Knu91]. customer [Pat92]. cutout [HF93]. cycle [MP90]. cycles [Fri92, Mur91b]. cylinder [enNAa94]. cylindrical [MS94b].

D [CXL94, Hei91, Jin94, LYX^{+}93]. damped [Jac90, UK94]. Darcy [BH93, HB93, HB91]. data [BP92, BCG91, Da93c, Da93d, Da93b, Eno90, IPP91, Koh92, SHR94, Vev92]. decision [DGP93, Mal93, Sar93, Yak92b, YLS93]. decomposable [Mal91]. decomposition [AR93, BS93, Cai93, Dos90, Hei93b, Yee93]. deficient [O’N90, SV94, VBH91]. defined [Fun91]. definite [Eva90a, Eva90b, ER94b, ER94a, Hua91, Sha94, Udw92b]. definition [BD94]. degree [Mur91a]. delay [EL91, FLZ91, FL94, GS94b, JV94, VB92]. delay-independent [FLZ91]. delays [Udw91, Udw92a]. demography [AR90a, AR90b]. dendrites [CFT^{+}93]. dependent [MSLB92, MS94c, PS94b]. Derivation [Jim94]. derivative [Sem92]. derivatives [Hus90]. Design [SS91, Bay93]. Detecting [Fri92, Mur91b]. determination [Mor90]. determining [BJQBQ94]. deterministic [Den90, Pat90]. develop [KS92a]. developing [RK93]. development [HB93]. devices [Ade94]. diagonal [Hua91]. diagram [NG94]. Dichotomy [ML90]. difference [BG94b, Dav94, Gyö90, Jim94, JF91, JNM92, JNF95, LG94, Lin94, MEZI93, NFJ94, Phi92, PPS94, Sel94, SA94, ZL90]. Differential [RBA92, Ada91, Ada93b, AR90b, Bai91, BKM94, CM90, Da93a, Da93b, EN94, ES92, ES93a, ES93b, ESS94, ES94, Ela94, EL91, GL92, Gro93, Heu90a, Heu90b, Jia94, JN91, JF91, Jód92, JL93, JV94, Kha91b, LS91a, LMS92, LWX92, Liu94, RK93, ŠNG92, Sha90a, Sha90b, SS94b, SS92, TT90a, TT90b, Unn90, VB92, XYY94, ZA91, ZXY94, AR90a]. differential-algebraic [Bai91, Kha91b]. diffusion [Asa92b, GS94a, HV90, KCK91, Yee93, Zas93]. dimension [JL93, Lap92]. dimensional [BH93, Bar91, BG94c, CS92, GL94, HV90, JM93, KCS90, Mur91b, NSV92, WM91]. dioxide [Ten90]. Dirac [Kha91a]. Direct [SH92, Knu91]. directed [PSMM93]. Dirichlet [MM94]. discontinuities [Dav94]. Discrete [MM91, Gol93, Jim94, KPP93, Kwo93, LAK94, Mal93, MDR94, PA94, Pat91, RTJL92, SS91, SW91, SY93, TLG90]. discrete-time [RTJL92, TLG90]. discrete-valued [SS91]. discretization [Da93c, Da93d]. disease [Yip91]. displacement [Kai91]. displacement-rank [Kai91]. distinguished [Pat92]. distributed [BJQBQ94, Den90, El 92, Udw91, Udw92a]. distributed-parameter [Udw91, Udw92a]. distribution [Sla93b]. Distributions [Pat92, MS93a, Sch91b]. divergences [PSMM93]. diversity [GG91a]. domain [BS93, Cai93, Dos90, GC94, Hei93b, SS91]. domains [Ham92]. dominance [Hua91, YLS93]. double [RRA90]. doubling [LAK94]. Drazin [Ji94]. droplet [SKCW94]. droplet-turbulence [SKCW94]. drug [LL94]. dusty [BH90, BH93, HB93]. Dynamic [SO90, Ben92, EB91, LS91b, SY93, SM94, UK94, Whi93]. dynamical [RW93]. Dynamics [LF94, BD90, BG94a, HL91, HLW94, Mos90].

Eckalbar [PBL92]. ecological [Den90].

economic [Avg94, KP91, Zhe93].

economical [NP92]. economics [KS92b]. economy [Pap92, Sos91]. Eddy [CXL94]. Editorial [Ano90b, Ano90e, Ano91a, Ano91b, Ano91d, Ano92b, Ano93b, Ano93d, Ano94e, Ano94g, Ano90a, Ano90c, Ano90d, Ano90f, Ano91c, Ano92c, Ano92d, Ano92e, Ano93a, Ano93c, Ano94a, Ano94b, Ano94c, Ano94d, Ano94f]. effect [Ham92, KK91, Kub94b, Raj94, ZXY94]. effectiveness [Ahm93]. effects [FE90a]. Efficient [How94, AH94, ES90, Gro93]. Eigenform [HLW94]. eigenvalue [Sha94, WJ92]. eigenvalues [Ben92, BG94c, How94]. elastic [enNAa94, Whi91]. electric [NA91]. electrocardiography [IM92]. electrochemistry [Mac93]. electromagnetic [LYX^{+}93]. element [Cha90a, CXL94, Da93a, Da93b, EN92, El 92, GC94, Hei93a, IBS94, Jac90, LYX^{+}93, Mei93, Siv92]. elimination [Moh94]. ellipsoid [NP92]. Elliptic [Son93, Ada91, Ada93b, Cas94, EHE93, KM94, LM94, LWX92, TH94, Vas91, Whi91]. elliptic-type [TH94]. elliptical [BJQBQ94]. embedded [HM94]. emerging [GKn91]. enclosure [Vol92]. enclosures [ZW90b]. End [Ram90]. endo [RR93]. energy [FV91, HJ92]. engineering [DGP93]. Engquist [Vul90]. Enhancements [CS92, Ada91]. entry [Ham94b]. environment [Pat90]. epidemic [ALMS90, Yak92b, Yip92]. epidemics [MO92, YGH90]. epidemiological [PR93]. Epstein [EHE93, TH94]. Equation [JV90, AS94a, AS94b, BJQBQ94, BG94b, Da93c, Da93d, EN94, FV91, FC93, Gol93, Gro93, GL94, HV90, Jac90, Jin94, Jód92, KN90, Kin93, MS94b, MEZI93, NSV92, PPS94, RC91, RK93, Sen93, SP94, Yee93, LAK94]. equations [Abd93, Ada91, Ada93b, AR90a, AR90b, Ayo93, Bai91, Cha90b, CZ92, CM90, Da93a, Da93b, ES92, ES93a, ES93b, ESS94, ES94, Ela94, EL90, Ge90b, GM94, GB90, GL92, GS94b, Gyö90, Heu90a, Heu90b, Jac92, Jia94, JL92, JN91, JF91, JL93, JV94, KT91, Kha91b, LS91a, LYC93, LG94, Lin94, LWX92, Liu94, Mit90, PS94a, Phi92, RBA92, RRA90, Sav93, ŠNG92, SR90, SU90a, SU90b, Sha90a, Sha90b, SA94, SS94b, Siv92, SS92, TT90a, TT90b, Unn90, VS90, VB92, XYY94, ZA91, ZL90, dK94]. equilibria [VCM91]. equilibrium [NP92, Ten90]. equivalent [LS93b]. Erratum [Cab95, JNF95, Sor97]. Error [GM94, Jac90, Jac92, AC94, CW92, CB94b, Da93b, GL94, Han92, HLW94, HS94, Jód92, MS92a, Mor90, Mor94, SH92]. error-free [Mor90, Mor94, SH92]. estimate [Han92]. estimates [AC94, Cai93, Da93b, GM94, HLW94, Jac90, Jac92]. Estimating [HS94, GG91a, RBR94, Sch91b]. Estimation [Whi91, Whi93, Ahm93, AB94, Bay93, Cic92, Cic93a, CH93, Cic93b, Faz91, IPP91, Sag91, Sla93a, TA93, Tes91c, TT90a, TT90b]. estimator [Jac93]. estimators [FS94]. Euclidean [Pes91]. eutrophic [GGR91]. Eutrophication [Gar90]. evaluating [ER94a, GC94]. Evaluation [Eva90b, Sar93, BG94a, Eva90a, ER94b, GM93, Hus90, Kha90, Kha92, Zhe93]. Evolution [KS92a, Den90, Ebe93, SU90a, SU90b]. Evolutionary [KCK91]. Exact [EHE93, JN91, FLYD94, Lan92, RTJL92, Yak92a]. exception [CJZ94]. Existence [GS94a, IST93, Koz91, KPP93, Gyö90, KP91, Pap92]. expansion [GL94, IST93, Liq91, MFS93, PS94a]. expectations [SY93]. experimental [Eno90]. experiments [SS91, Yip93]. explicit [Sim92]. Exponential [ADM91, Kum92a, MS93a]. exponential-polynomial [MS93a]. exponentially [Sel94]. exponentials [MMP90]. expression [Ji94]. extended [CB94a, CB94b, Knu91]. extension [Pat90]. extraction [Bha94]. extrapolation [BG94c, IST93]. Extremal [CN90]. Extremum [Sor94, Sor97].

facility [She90]. factor [GG91a]. factors [SS91, TCS92]. Factory [CDW91]. facultative [Zag91]. fail [SH92]. fail-proof [SH92]. families [MS93a]. family [Jim94]. Fast [Mor90, Mur91b]. fatal [Yip91]. fate [Pat90]. feedback [TFM93]. Fermi [Kha91a]. field [HM94, VWSS94]. fields [Bar91, LYX^{+}93, SW91]. filled [Ge90a, GQ90]. filter [CB94a, CB94b]. filtering [Heu90a, Heu90b]. final [PY94]. finance [Sha90a, Sha90b]. find [Bar91]. Finding [Ge90b, Sav93, XYY94, ZG90, Jia94, PP93]. Finite [Da93a, El 92, Hei93a, Lin94, Mei93, Pic92, Cha90a, CXL94, Da93b, EN92, EL91, FG94b, IBS94, Jac90, Jim94, LYX^{+}93, MEZI93, RC91, Sel94, Siv92]. finite-difference [MEZI93]. finite-element [Jac90, Siv92]. Finite-horizon [Pic92]. First [And93, Abd93, Ayo93, BKM94, Cab94, Cab95]. First-Order [And93, Ayo93, Cab94, Cab95]. fitted [Sel94]. fitting [BCG91, BG94c]. FitzHugh [Jac92]. fixed [Gun94, HF93, Pes91]. flat [HM94]. fleet [MFS93]. flip [KR91]. flow [BH93, Ezz94, Faz91, HB93, Ham92, Ham94b, Han92, Mei93, Mei94, MS94a, NG94, Raj94, RWB94, Uen91, VWSS94, WN91, Zhu94]. flows [BTW94, KS93a, Raj92]. Fluctuon [Con93c]. fluid [BH93, HB93, Han92, RWB94]. Flux [Dav94]. focused [Lan92]. fog [BD94]. food [LF94, SPD93, Zag91]. food-chain [LF94]. force [Con93c]. forced [Ela94]. forest [TDK94]. form [Gun92, JNM92, NFJ94]. formal [Pat93]. formalism [Kon92]. Forms [Gun93, Mit92]. formula [Kha91a, ZXY94, HS94]. formulas [CWL92, CS93]. Formulation [RW93, BTW94, Cha90a, Cha90b, LS93b, MS94a]. forward [Lan92]. forward-scattering [Lan92]. foundations [Wer93]. founding [Mat93]. Four [JV90, Eva90b]. Fourier [RWB94, RRA90, dFFG94]. fourth [Gup90, GK94]. fourth-order [Gup90, GK94]. fractal [Yan94]. fractional [ES92, ES93a]. Fredholm [GB90, KN90, RRA90, Sen93]. free [BTW94, Ezz94, Mat93, Mor90, Mor94, Raj94, SH92]. freedom [Mur91a]. frequency [SS91]. frequency-domain [SS91]. friend [Cas91]. Frobenius [DDL93]. fueling [CK90]. Function [ZG90, Ayo93, BG93, BG94c, Car90, Ge90a, GT91, KCS90, LW90, Mac93, Mal91, RC91, RTJL92, Sor93, Sor94, Sor97]. functional [ES94, GT91, GL92, LYC93, Pic92, VCM91]. functions [AaHM93, BP92, GQ90, Ioa92, Kha92, Kin91, Koł94, LMV91, MDR94, MW93, Son93, Sri94, ZM91, ZW90a]. fusion [CK90]. Fuzzy [BBD91, BD94].

Galerkin [Asa92a, Jac92, MSLB92, ZM91]. gas [BH90]. Gaussian [Lan92, MS92a]. Gel’Fand [Kin93]. General [ES90, NFJ94, RK93, TA93]. Generalizations [ZED91]. Generalized [PP93, Arg90, Arg94, BP92, EHE93, Gol93, Mal91, Sen90, Sri94, TH94]. generated [How94]. generating [Sri94]. Generation [SW91, CS92, GSK93, Ham94a, KM94, Knu91, Knu93, SW92, SYS94, Son93, Spi90, YS94]. generator [And90]. Generic [ARM94]. geometric [CWL92, Liq91, MES92]. Ginzburg [Kwo93]. Global [CUK92, LS91b, PPS94, Sor93, Ge90a, GQ90, Moh90, SW94, Sla93a]. globally [GQ90]. Gordon [FV91, JV90].

Goursat [Waz93]. gradient

[AH94, HV90, Jac93]. gradostat [Zag92]. Graduating [BP92]. Gram [TM92]. gramicidin [Kub94b]. granule [CFT^{+}93]. Green [Ayo93, MS94b]. Greenberg [Mad90]. Grid [And90, CS92, Ham94a, KM94, Knu91, Knu93, KCS90, SYS94, Son93, Spi90, YS94]. grids [Ada93b, SW92]. Gronwall [ZED91]. group [MES92]. growth [Avg94, Chu92, Chu94, KP91, KPP93, Pap94, Ten91].

half [Reg94]. half-wave [Reg94]. Halley [CAQ93]. Hamiltonian [Fri92]. Hankel [Kha92]. Hansen [Mad90]. harmonic [Jac93]. health [BBD91]. health-risk [BBD91]. heat [HM94, Jin94]. hemisphere [MS94b]. hepatic [Bha94]. heredity [FC93]. Hermite [Kha91a, LMV91]. Hermite-type [LMV91]. hierarchical [DHK90, Mal93]. High [DDL93, VWSS94]. Higher [Abd93, LW94, IST93, JNM92, JNF95, NFJ94, ŠNG92]. higher-order [JNM92]. highly [MEZI93]. Hilbert [Cro93, IPP91]. Hilliard [Fre90]. HIV [LL94]. holding [PBL92]. hollow [enNAa94]. homogeneous [IPP91, VB92, Wam92]. homotopy [KT91, LqGq93, MS92c, XYY94]. Hooke [HLW94]. horizon [Avg94, KP91, Pap94, Pic92]. Householder [Cic92]. hp [IBS94]. hp-finite [IBS94]. Hubbell [EHE93, TH94]. Hybrid [Swo91]. hydrogen [CK90]. hydrology [SU90a, SU90b]. hyperbolic [Ayo93, CS92, LYC93, Li94]. hypergeometric [Yip93]. Hypernumbers [Mus94].

identification [GSK93, KKU91, Sag91]. II [BH93, Cha92d, Da93d, ES93b, Eno90, MK93]. III [Cha92a, ER94b]. imaging [FE90a]. imbedding [Ayo93, Sag91]. Implantation [Ada93a]. implementation [BS93]. implementations [SH92].

Implementing [Bre91, KMR^{+}91]. implicit [Abd92, HL91, JF91, JNF95, Mit90, NFJ94]. importance [YLS93]. imprecisely [MDR94]. improvements [Mad90, Vul90]. impulse [ZXY94]. Impulsive [LMS92, BKM94, EL90, SS92]. income [Chu94]. incomplete [GSK93, SO90]. incompleteness [Cha92b, Gun94]. incompressible [BH90]. increase [JL93]. independence [CJZ94]. independent [FLZ91, Jol92, SA94]. Index [Ano90g, Ano91e, Ano92f, Ano93e, Ano94h]. induced [KS93a, MM94]. industry [SPD93]. inequalities [BKM94, GT91, ZED91]. inertias [Udw92a]. Inference [HT90a, HT90b, Yip91, Yip92, Yip93]. infinite [Abd92, Avg94, BG94c, HM94, Kha90, KP91, Pap94]. infinite-dimensional [BG94c]. infinity [Sim92]. Information [Cha92b, Ebe93, GSK93, MGL92, SO90]. Information-theoretic [Cha92b]. inhomogeneous [Asa92b]. initial [Da93c, Da93d, Da93b, Gyö90, JNM92, Lin94, Sim92, Sim94, SV92a, SV92b, SP93]. initial-value [Sim92, Sim94]. input [Bay93]. inscribed [NP92]. instabilities [TW90]. insulin [Bha94]. integer [Mor94]. integral [EHE93, GB90, GM93, GL94, Gup90, RRA90, SR90, Sen93, TH94, dK94]. integral-type [Gup90]. integrals [Eva90a, Eva90b, ER94b, ER94a, GC94, Kha90, Kha91a]. integrated [AaHM93]. integration [Koh92, Red90a, Red91, Sim92, Sim94]. integro [Da93a, Gro93]. integro-differential [Da93a, Gro93]. integrodifferential [Abd93, Da93c, Da93d, EL90, NSV92]. interacting [Pat91, Zag91]. interaction [HM94]. interactions [SKCW94, SP93]. Interactive [SYS94, TDK94, AaHM93]. interface [Mus94]. intergrodifferential [Rua92]. internal [Mos90, Sla91].

International [Sla93a]. interpolation [ADM91, CWL92, LMV91, Ram90]. interpolations [Wan92]. interpolatory [Beh92]. interpretation [Mos90]. Interval [Mad90, Moh94, Kop90, LS91a, SW94, SS92, ZW90b]. intervals [RC91]. intravenous [LL94]. Introduction [Tes91a, Unn90]. Invariant [Kop90, Sag91, Ayo93, DGV94, Gyö90, Sen90]. inverse [BL91, IM92, JM93, Ji94, Kag91b, Pen93, RBR94, Sen90]. inversion [FLZ92, Wan93]. invert [LSS94, LSVS94]. invertibility [MR91]. Invertible [DGV94]. inverting [Kwo93]. investment [Hoj94, ZP92]. involving [Avg94, MFF92]. ion [Kub94b]. IRK [Abd92]. irrational [Eva90a, Eva90b, ER94b, ER94a]. Irregular [Mur94]. irrigation [MN91]. issue [DGP93, SS94b, Unn90]. iterate [BG94c, Gol93]. iterated [GL94]. iteration [NAM93, SB92]. iterations [LWX92]. Iterative [CZ92, JL92, MES94, EL91, Faz91, Udw92b, UK94]. IV [Cha92e, ER94a].

jackknife [Bre91]. Jacobi [CS93, EA94]. jet [CJF93]. joined [MS94b]. JOR [Udw92b]. Jordan [Sen90]. July [Cab95].

Kalaba

[Cas91, KKU91, Kag91b, Tes91b, Tes91c]. Kalman [CB94a, CB94b]. kernel [Da93c, Da93d, NSV92]. kernels [Mor90]. kidney [BL91]. kind [Abd93]. kinetics [Ten91]. kinship [Sch91a]. Klein [JV90]. knapsack [Dro90]. knot [Beh92]. Kornai [Sos91]. Kronecker [MFF92]. KSR [Zhu94]. KSR-1 [Zhu94]. Kutta [Abd92, HS94, And93].

lag [Sim92, Sim94]. Laguerre [RS92]. laminar [Han92]. Landau [Kwo93]. Laplacian [HB91]. Lapwood [BH93, HB93, HB91, HB91]. large [She90]. largest [BG94c]. lattice [Mur91a]. law [HLW94]. laws [Con93a, Jim94, Li94]. Lawvere [Gun94]. layer [AS94a, AS94b, HM94, Raj94]. leader [SMO91]. Learning [GSK93, SM94, AH94, KCK91]. Least [LYX^{+}93, CXL94, TM92]. Least-square [LYX^{+}93]. least-squares [CXL94]. Leibniz [Mur92]. length [Faz91, KK91, Sch92]. Leslie [MR91]. Level [Flå92, RWB94]. Level-constrained [Flå92]. levels [CB94b]. Levison [Mor94]. Levitan [Kin93]. LH [KMR^{+}91]. library [Hus90]. life [Con93c, GKn91, Gun92, Gun93, Gun94, KG91]. like [Arg94]. Limit [Koł94, CM90, Ji94, Mur91b]. limitations [Pat93]. limited [Lap92]. limiters [Dav94]. limiting [Zag92]. line [Spi90]. line-spacing [Spi90]. Linear [ES93a, KKM90, Reg94, Cab94, Cab95, CH93, Da93b, FLZ91, Ge91, Hoj94, Kop90, Lop92, MS94c, SR90, SS94a, Yak92a, ZA91]. linearized [Uen91]. linkage [MS94c]. Liouville [WJ92]. lipid [KK90, Kub94a]. LISP [Cha92c, Cha92d, Cha92a, Cha92e]. living [Kon92, MS93c, MK93]. local [Koč92]. localized [ALMS90]. locating [DFM94]. location [Pes91, She90]. logarithmically [GM93]. logic [MO93, TGK91]. logistic [FC93]. loop [Mur91a]. loss [MN91]. low [KCS90, RWB94]. lower [OT92, OT93].

Mach [RWB94]. macroscopic [MO92]. magnetic [HM94]. magneto [enNAa94]. magneto-elastic [enNAa94]. magnetohydrodynamic [Mei93]. magnetohydrodynamics [Mei94]. magnification [Mit90]. magnitude [LMV91]. making [DGP93, Sar93, YLS93]. management [TDK94]. managerial [Sar93]. Manifold [Beh90]. mappings [MM94]. maps [KR91, Szi90]. market [Mos90, Sos91, SM94]. Markov [LS93b, Pat90, Pat92]. Markov-renewal [Pat90, Pat92]. Markovian [MES92]. martingale [Yip92]. massive [MO93]. matching [AR93]. Material [HKLS93, HL91, HLW94, Whi93]. materials [HLW94, PZ94]. Math [Cab95]. Mathematical [Chu92, KG91, ALMS90, AR90a, AR90b, CFT^{+}93, Mur94, Ten90, Zhe93]. Mathematics [Ano92a, Kag91b]. Matrices [MS94b, How94, Hua91, LSVS94, Udw92b]. Matrix [MES92, BG94c, CC94, Cha90b, FLZ92, JN91, Jód92, JMV93, JL93, LSS94, Sha94, SP94, TCS92, Wan93]. Matrix-geometric [MES92]. maximal [LS91a, Pap92]. maximum [Cha93]. Maxwell [Cha90b]. MCDM [MTP93]. mean [Bha94, FLG92]. means [SV94]. measles [ALMS90]. measure [Lap92, ZXY94]. measured [Vev92]. measurement [Pat93]. measures [Kop90, MS93b, MGL92]. mechanics [YF91]. mechanism [WM91]. media [BH90, BH93, HB93, Zhu94]. median [FS94]. medical [FE90a]. medicine [FEGL90, FE90b]. medium [Ezz94, HM94, Ten91]. membrane [KK90, KK91]. membranes [BMS92, Kub94a]. memories [KKU91, KKM90]. memory [El 92]. mesh [KB93]. meshes [WN91]. Method [And93, HKLS93, ZG90, Arg90, AC94, Asa92a, Asa92b, Bai91, BJQBQ94, BS93, Cab94, Cab95, Cha90a, CWL92, CAQ93, Cro93, Da93b, EN92, Faz90, Faz91, GC94, Gro93, Hei91, JR93, Jia94, Jin94, JL92, JF91, KS93a, KS93b, Knu91, Koč92, Koh92, KK91, LYX^{+}93, Li94, Mad90, MSLB92, Mit90, Moh94, O’N90, OT92, OT93, PS94a, PP93, PZ94, Red90a, Red91, Sch91a, Sel94, Sim92, Spi90, Ste94, TCS92, Udw92b, Vev92, XYY94, Yee93, ZM91]. methodology [Sav93, Yak92b]. methods [Abd92, Abd93, Arg94, Ben92, CW91, CZ92, Cic92, Cic93b, Da93a, Dos90, EL91, GM94, GB90, Hei93a, Hei93b, HV90, Hua91, IST93, JMV93, KB93, LqGq93, NP92, SA94, Sim94, UK94, WM91, ZW92]. MHD [MS94a]. Microelectronics [SP94]. Microscopic [MO92]. midpoint [AC94]. migration [Sla91, Sla93a]. mild [KN90]. Minimal [Sha94, HJ92, LS91a, Mit92, Sim94]. minimal-energy [HJ92]. minimax [Ge91]. minimization [VCM91]. minimizing [BD90]. minimum [Sor93]. minimums [BG93]. Mixed [HM94, Cha90a, JL92]. Model [Gar90, BL91, BG94b, CFT^{+}93, Fre90, GP94, IM92, KPP93, KK90, Kub94a, LL94, Mos90, Pap94, Pat90, PS94b, PR93, PBL92, Sos91, SHR94, SMO91, Ten90, Whi91, Whi93, Yak92b, Yip92, Yip93, Zhe93, Con93c]. modeled [Szi90]. Modeling [BMS92, YF91, Bog91, Han92, Kag91b, Ros93, YGH90]. models [AS90, AS92, AB94, BH93, Ben92, EB91, HT90a, HT90b, HB93, HB91, Ham94b, KM94, MN91, NP92, Pat93, Pat91, Raj92, Ros93, RW93, Sla91, SO90, TA93, VS90, WM91]. modified [BTW94, KMR^{+}91, LS91b, TM92]. modifier [KK91]. Modular [Car90]. modulated [Reg94]. molecules [KK91]. monitoring [NA91]. Monotone [EL91, LWX92, Vol92, AS92, Cab94, Cab95, JL92]. Monte [RBR94]. MOSI [Bar91]. motion [BH90, KU92, RW93]. MUDPACK [Ada91, Ada93b]. MUDPACK-2 [Ada93b]. Multi [YLS93, MDR94, Wam92]. Multi-attribute [YLS93, MDR94]. multi-homogeneous [Wam92]. multiattribute [Mal91, MW93]. multiblock [SW92]. Multicriteria [MFS93, Tes91c]. multicriterion [CDW91, DGP93]. multidelay [FLZ91]. multidimensional [Kon92]. multidisciplinary [SHR94]. multifacility [Pes91]. multigrid [Ada91, Ada93b, CWL92, Hei91, Hei93a, Hei93b, HV90, Hua91, Koč92, MP90, NW90, Spi90, Vol92, Zas93]. multigroup [Zas93].

Multilevel [Vas91]. multiobjective [AaHM93, TDK94]. multiparameter [ASZ94]. multiphase [Zhu94]. multiple [KS92a, Mal93, Sav93, WN91]. multiple-criteria [Mal93]. multiplicity [HSZ94]. multisector [Pap92]. multiserver [MES92]. multisinusoidal [Bay93]. multisteps [JMV93]. multivalued [TGK91]. multivalued-logic [TGK91]. mutual [MGL92]. mutualism [Zag91].

Nagumo [Jac92]. narrow [RW93].

Naslund [Sar93]. nations [Chu92, Chu94]. Natural [MS93a]. naturality [NAM93]. Navier [VWSS94]. Necessary [FLZ91, BD90, Gun92]. neighborhood [Mit90]. network [FLZ92, FLYD94, Pat92, PS94b, Wan93]. networks [AH94]. Neumann [HKLS93, HKLS93]. Neural [FLZ92, FLYD94, AH94, Wan93]. neuron [CFT^{+}93]. neuronal [HT90a, HT90b]. neurons [KCK91]. neutral [CUK92, GL92, GS94b, LG94]. neutron [Zas93]. Newton [Arg94, Mad90, Mur92, ZW90a, ZW92]. Newton-like [Arg94]. Newton-type [ZW92]. Nicolson [Jin94]. no [Ham92]. no-slip [Ham92]. node [Pat92]. noise [TW90, VS90]. noise-in-variables [VS90]. Non [Da93b]. Non-smooth [Da93b]. nonclassically [UK94]. Noncollocated [Udw91, Udw92a]. nonconvergence [SB92]. nondifferentiable [SW94]. nondispersive [Udw91, Udw92a]. nonequidistant [Vul90]. nonhomogeneous [NFJ94]. noniterative [Faz90]. Nonlinear [ARM94, Gar90, JV90, SA94, Abd92, Asa92b, CN90, Cab94, Cab95, CZ92, CUK92, Da93a, ES93b, EL90, FLG92, Ge90b, GP94, GT91, GS94b, Gup90, HSZ94, KT91, Kum92a, LqGq93, LWX92, MM91, MW93, Pen93, PPS94, Sav93, ŠNG92, Ste94, TA93, TFM93, Vev92, Vol92, Whi91, Whi93, YF91, PS94a]. nonlinearities [KS92a]. Nonlinearly [TLG90]. nonnegative [Gyö90]. nonoscillation [ES93b]. Nonparametric [Ahm93, Sch91b, PY94, TT90a, TT90b]. nonsmooth [Da93d]. nonstandard [Kon92, SV92a, SV92b]. nonstationary [Szi90]. nonsymmetric [Jód92]. nontrivial [Gyö90]. Norm [Cai93, Cro93]. normal [BG94c]. norms [Da93b]. not-a-knot [Beh92]. note [AS94a, Cha93, CAQ93, FG94a, Mac93, MS92a, Red91, Sem92, Sen93]. nuclear [FEGL90, FE90b]. nuclear-medicine [FEGL90]. null [CUK92]. null-controllability [CUK92]. number [Cha93, CB94a, KCS90, RWB94, Wam92]. Numerical [CJF93, IPP91, JV90, JMV93, KR91, LS93a, Red90b, SKCW94, WM91, ADM91, Asa92b, FV91, GM93, Ham94a, JF91, KN90, Kha90, Kha91b, Kha92, Koh92, LW90, Li94, LWX92, LF94, PS94b, Red90a, Red91, Sim92, Sim94, SV92a, Waz93]. numerical-integration [Red91]. NURBS [SYS94].

observation [Cic93b, CB94b]. observations [CB94a, CB94b, FC93]. observers [CB94a]. obsevability [MM91]. obstacles [LSP91]. ODE [Liq91]. ODEs [And93, IST93, SB92]. oligopolies [SY93]. oligopoly [SO90]. one [Kum92a]. operation [EB91, PL93]. operational [CS93]. operator [CZ92, DDL93, JL92, Kwo93]. operators [Cai93]. optical [FE90b]. optics [Pen93]. Optimal [Avg94, Chu94, CH93, Gar90, TT90a, TT90b, GGR91, KP91, Pic92, RTJL92, TLG90]. optimal-control [TLG90]. Optimality [Cha91]. optimization [AaHM93, Ge90a, GQ90, Kag91b, Moh90, SW94]. Optimum [Kum92a, Hoj94, ZP92]. orbits [DFM94, FG94b]. Order [And93, LW94, Abd93, Ayo93, BKM94, CN90, Cab94, Cab95, DDL93, Ebe93, EN94, ES93b, ESS94, ES94, Ela94, FT94, GM94, GL92, GS94b, Gup90, GK94, IST93, JNM92, JL93, JV94, JNF95, KB93, KS93b, LG94, Lop92, NFJ94, OT92, OT93, ŠNG92, Sim92, Sim94, SV92b, YLS93, ZA91]. orders [ES93a, Ioa92, MTP93]. ordinary [Jia94, ŠNG92, XYY94]. ORESTE [MTP93]. Organ [FEGL90]. origin [Kai91]. orthogonal [Kin91]. Orthogonality [Sri94, RS92]. orthonormal [RK93]. orthotropic [enNAa94]. Oscillation [ES93b, GL92, LYC93, LG94, KS92b]. Oscillations [Mur93, Phi92, Mur94, ZA91]. Oscillators [ARM94]. Oscillatory [Ela94, ESS94, ES94, GS94b, Kha90]. Osher [Vul90]. Ostrowski [PP93]. other [Mat93]. outflow [Sem92]. overlapping [WN91].

Pacific [Sim92]. package [Ada91]. Padé [Kum92b]. Parabolic [SP93, CW91, CW92, Da93c, Da93d, Da93b, Kor92, Lin94]. paradox [CM90]. Parallel [HL91, BS93, Cro93, EA94, Koł94, MO93, SH92, Zhu94]. parallel-series [Koł94]. parameter [Cic93a, MSLB92, MS92b, SA94, Siv92, Udw91, Udw92a]. parameter-dependent [MSLB92]. parameters [BJQBQ94, Lan92, Whi91, Whi93, ZXY94]. paribus [Mat93]. part [BH93, HB93, MS93c, MK93]. partial [Ada91, Ada93b, BKM94, BG94b, Da93b, JF91, LWX92]. partition [SB92]. partitioning [Cha91]. parts [Car90]. patching [Hei93b]. paths [Mur91a]. Pattern [Wu94]. PDE [Fun91, KM94]. PDEs [Vol92]. pellet [CK90]. penalty [RTJL92, Yak92a]. Perfect [AB94]. performance [Ahm93]. Periodic [ŠNG92, CN90, CW91, CW92, Jia94, Kor92, LAK94, Liq91, Liu94, Mor90, OT92, WJ92, XYY94]. Perron [DDL93]. persistence [Zag91]. person [SMO91]. Perturbation [Red90b, BS93, JR93, Red90a, Red91, Sel94, Sem92]. perturbed [Bai91, KB93, Lop92, VB92]. phase [Ham94b, Sim92, Sim94]. phase-lag [Sim92]. phased [Bay93]. phenomena [PA94]. phenomenon [FE90a]. photometric [Koz91]. Physical [Sla91, Mus94]. Physical-based [Sla91]. physics [Mat93, Wer93]. piecewise [Beh92, Kop90]. piezo [Bal94]. piezoelectric [Ade94]. pipe [Mei93]. plane [Knu93]. planning [KP91, Rap92]. plant [Bay93, RBR94]. plasma [CK90]. plasma-arc [CK90]. plate [HM94, RW93, Whi91, Whi93]. plates [Bal94]. point [Faz90, GM94, Gun94, KS93b, LMS92, Lop92, ML90, MFF92, Pes91, SV94, Sor93, Szi90, TM92, Udw91, ZW90a]. point-to-set [Szi90]. Points [LW94, FT94, JL93, Lop92, Sor94, Sor97]. pointwise [Pic92]. Policies [Gar90]. Polynomial [Rug93, Beh92, HV90, MMP90, MS93a, MS93b, MS92c, MS92b, VBH91, Wam92]. polynomials [CS93, LS93a, MS93b, MM94, RS92, Sri94, TCS92, dFFG94]. Population [BD90, ES90, Sla93b]. populations [LF94, Pat91, PS94b, Zag91]. porous [BH90, BH93, Ezz94, HM94, HB93, Ham94b, Zhu94]. position [Bal94]. positive [Hua91, Sha94, Udw92b]. positive-definite [Udw92b]. positone [ASZ94]. possible [Gun92, Mos90]. potential [CFT^{+}93, MS94b, WN91, Wer93]. power [DHK90, NA91]. Practical [FL94, FLG92, SHR94]. prebiotic [Ten91]. preconditioned [HV90]. preconditioners [Vas91]. preconditioning [Dos90, Hei93a, Rug93]. prediction [Zhe93]. Preface [DGP93]. presence [TW90]. presentation [CS93]. pressure [Bar91, Cha90a, VWSS94]. pressure-correction [Bar91]. price [GP94]. prices [KPP93]. Prime [KCS90]. principle [Mos90, Mur91b]. principles [FE90b]. prior [Sch91b]. priori [CH93]. probabilistic [BBD91]. Probability [MS93b, MS93a]. problem [Asa92b, BS93, BL91, Cha90a, CXL94, EN92, Gup90, GK94, IM92, IPP91, JL93, KP91, LM94, Moh94, Pen93, Pes91, Sim94, Uen91, War94, Waz93, ZXY94]. Problems [Red90b, ASZ94, Asa92a, Avg94, Bai91, CN90, Cab94, Cab95, Cas94, CW91, CW92, EL90, Faz90, GS94a, Ge91, GM94, HSZ94, Hei93b, JR93, JNM92, KB93, Kag91b, KS93b, KS92a, Koč92, LMS92, Mal93, MSLB92, ML90, MFF92, OT92, OT93, PA94, Red90a, Red91, RTJL92, SV94, ŠNG92, Sel94, Sem92, SS94a, She90, Sim92, SV92a, SV92b, Ste94, TLG90, TFM93, TM92, Vas91, WJ92]. procedural [Mad90]. procedure [BTW94, GG91a, Siv92, Yip92, Yip93]. process [TM92]. processes [CC94, GSK93]. producer [SM94]. product [Cai93, Mit92]. products [MTP93, MFF92]. program [Cha92c, Cha92d, Cha92a, Cha92e, PZ94, Zhe93]. program-size [Cha92c, Cha92d, Cha92a, Cha92e]. programming [EB91, Flå92, Ge91, Hoj94, KKM90, TDK94]. programs [Pap92, Yak92a]. project [CDW91]. projection [Cro93, GB90]. projections [Kha94]. projective [Yan94]. Projector [Dos90]. proof [Gun94, SH92]. propagation [enNAa94]. properties [Liq91, MM94, PSMM93]. property [Jac90, NG94, RS92]. pruning [LSS94]. Pseudospectral [Fun91]. Ptâk [AC94]. putting [PZ94].

Quadratically [SV92a]. quadrature [MS92a]. quadrilateral [Wan94]. Qualitative [GP94, LL94]. quality [Bog91, SPD93]. Quantitative [Han92]. quartic [SV94]. quasi [JM93, JL92]. quasi-solutions [JL92]. quasi-three-dimensional [JM93]. quasiconcave [AaHM93]. quasiconvex [AaHM93]. quasisteady [Raj92]. queue [MES92]. queueing [Sos91]. queues [MES94]. quintic [Ste94].

racemic [Ten91]. radial [Cas94]. Radiation [Rap92, FE90a, HM94]. railgun [CK90]. Randles [Mac93]. Random [Ben92, SU90a, SU90b, LS93a, PY94, SW91]. randomly [How94]. rank [Kai91]. Rapid [VCM91]. rate [Bal94, Cic93b]. rates [CB94b]. rather [Mat93]. Rational [Jód92, O’N90, PZ94]. reaction [KCK91, Yee93]. reaction-convection-diffusion [Yee93]. reaction-diffusion [KCK91]. reactor [Zas93]. reactors [CK90, Faz91]. real [Sha94, Wan93]. real-time [Wan93]. realistic [AB94]. realization [LS93b]. recapture [Yip93]. reclamation [GGR91]. recognition [Wu94]. reconciliation [Vev92]. Reconstruction [FG94b, Kha94]. reconstructions [Li94]. recourse [Yak92a]. recourse-constrained [Yak92a]. rectification [Reg94]. recurrent [Wan93]. recursive [NAM93]. reduction [Jac93, Sen93]. reductionism [Con93a]. reflection [NSV92]. regard [BD90]. regeneration [BD90]. regions [SW91]. Regression [Pat91, Kum92a, TA93, VS90]. regular [JNM92, JL93]. regularization [dFFG94]. regulator [TFM93]. Reid [ZED91]. related [CS93, Udw92b]. relations [Sri94]. Reliability [NA91, Koł94, PL93]. reliability-based [PL93]. reliablity [AB94]. relief [VWSS94]. Remarks [Cab95, Kai91, OT92, OT93]. renewal [Pat90, Pat92]. representation [ES90, Mor90]. representations [Sla93b]. reservoir [EB91]. residue [Eva90a, Eva90b, EHE93, ER94b, ER94a]. Resolution [Lap92, Ada93b, Dav94, FG94b, FE90a, Kha91b, Sch92]. Resolution-limited [Lap92]. resonance [GK94]. resources [PL93, TDK94]. respect [YLS93]. results [HSZ94, Udw92b, ZL90, dFFG94]. retarded [EN94]. reverse [NG94]. review [AR90a, AR90b, Ham94b, MU90a, MU90b]. revisited [Son93]. revolution [Wer93]. Riccati [JN91, Jód92, LAK94]. Richtmyer [HKLS93]. ridge [Cic93a, Cic93b, Cic92]. ridge-type [Cic93a, Cic93b, Cic92]. Riemann [Ayo93, IPP91]. risk [BBD91]. river [Bog91]. RLS [FLYD94]. Robert [Tes91b, Tes91c]. Robust [TA93]. root [PP93]. root-finding [PP93]. Roots [ZG90, Sav93]. Ross [FG94a]. rotation [VCM91]. Routing [MW93]. row [LSS94]. row-pruning [LSS94]. Rules [Con93a]. Runge [Abd92, And93, HS94].

s [ASZ94]. s-shaped [ASZ94]. safety [VWSS94]. Saffman [BH93, HB93]. salesman [Dro90, War91, War94]. sample [Bha94, BP92, CC94]. sampling [Cic93b, MGL92, PSMM93, RBR94]. scale [She90]. scattering [Lan92]. Scheme [HKLS93, CS92, Jol92, KCS90, Lop92, MEZI93, Vul90]. Schemes [JV90, FV91, HL91, Jim94, Sel94, SV92a]. Schmidt [TM92]. Schroeder [Bay93]. Schroeder-phased [Bay93]. Screening [MDR94]. search [Bad93, Cha91]. secant [Arg90]. second [CN90, ES93b, ESS94, ES94, Ela94, FT94, GM94, GL92, JL93, JV94, KS93b, LG94, Lop92, OT92, OT93, Sim92, Sim94, ZA91]. second- [OT92, OT93]. second-order [ES93b, GM94, JL93, JV94, Sim92, Sim94]. section [Bog91]. sectional [Cha91]. sectionally [Ioa92]. seek [Hoj94, ZP92]. selection [CDW91, TW90]. Self [KS92b, BG94a, MSLB92, Shi93]. self-adjoint [MSLB92]. self-creation [Shi93]. self-evaluation [BG94a]. Self-sustained [KS92b]. semi [Abd92, HM94, WJ92]. semi-infinite [Abd92, HM94]. semi-periodic [WJ92]. semidiscrete [Jac90, Jac92]. semigroups [Opp92]. semiinfinite [RC91]. semilinear [CW91, CW92]. semimartingale [TT90a, TT90b]. semipositone [KS92a]. Sensitivity [Cic93b, CB94a, CB94b, Eno90, Pap94, Tes91b, dK94]. sensors [Bal94]. Separated [And93]. separation [Mur91b]. series [BCG91, BG94b, EHE93, Kin94, Koł94, Kum92b, RRA90, Wu94, YF91]. series-parallel [Koł94]. servers [MES92]. services [Pat92]. Set [Sch91a, Bad93, BBD91, Bay93, Kin91, Szi90]. Set-spectrum [Sch91a]. sets [RK93, Yan94]. Sevcik [Mac93]. several [AB94, SA94]. shallow [HF93]. Shape [Bal94]. shaped [ASZ94]. Sharp [Gyö90]. shells [HF93]. shift [DGV94]. shift-invariant [DGV94]. shocks [HLW94]. shooting [GM94, WJ92]. shortage [Sos91]. shortest [Wai90]. sign [Gup90]. significance [HB91]. simple [KKM90, Yip92]. SIMPLER [Bar91]. Simulating [Zhu94]. Simulation [Ade94, Raj92, ALMS90, Raj94, WN91]. simulations [LF94, RBR94, SS94b]. simultaneous [VS90]. simultaneous-equations [VS90]. Sinc [MSLB92]. sine [Eva90b, FV91]. Sine-Gordon [FV91]. Single [Ham94b, Li94]. Single-phase [Ham94b]. Singular [Red90b, BS93, CW91, CW92, Da93c, Da93d, EN92, GB90, GM93, Gro93, JR93, JNM92, JL93, LM94, RBA92, Red90a, Red91, Sel94, Sem92, SR90, Sen93, ZXY94]. singularity [Ioa92, KN90]. singularly [KB93, Lop92, VB92]. singularly-perturbed [KB93]. sink [BH93]. sinusoids [Reg94]. size [Cha92c, Cha92d, Cha92a, Cha92e]. slackness [Moh90]. slip [Ham92]. Sloan [Gol93]. Slope [ZW92, ZW90b]. slotted [Sch92]. slow [Mur91b]. slow-fast [Mur91b]. slowly [SS94a]. small [Bha94]. smallest [BG94c]. smooth [Da93c, Da93b]. smoothed [Sla93b]. smoothing [Jac90, Sla93a]. software [Ada91, Ada93b, MTP93]. solid [Abd92, CK90]. solid-hydrogen-pellet [CK90]. solidification [Fre90]. solitons [Wer93]. Solution [ARM94, Cha90b, MFF92, RC91, EHE93, Gro93, Jia94, JN91, JF91, JNM92, KN90, Kin93, Kin94, Liq91, MES92, MES94, NFJ94, SR90, SS94a, Siv92, Waz93, Yee93]. Solutions [RRA90, SV92b, AS94b, AR93, CN90, Cas94, CW91, CW92, ESS94, ES94, Ela94, Ge90b, GS94b, JL92, Jód92, JV94, LS91a, Lin94, LWX92, Liu94, MS92c, OT92, OT93, RK93, She90, TM92, XYY94, Zag92]. solvability [GK94]. solve [Mor94]. solvents [TCS92]. solver [NW90]. Solving [Ge91, JL93, KT91, SV94, Abd93, GT91, HV90, IST93, JR93, KS93b, LqGq93, Mit90, NP92, PS94a, Pes91, Red90a, Red91, Sel94, VBH91]. Some [FC93, HSZ94, MM94, Sim94, Udw92b, Vul90, FS94, GB90, IST93, Koł94, Mad90, TT90a, TT90b]. space [Ezz94, Liu94]. spaced [Koh92]. spaces [Arg90, AC94, CAQ93, Cro93, KCS90]. spacing [Spi90]. spanning [Moh94]. spatial [FE90a]. Spatially [Den90]. Special [SS94b, War94, DGP93, Sim94, SP94, Unn90]. specifications [EB91, NAM93]. Spectral [Hei93b, Hei91, Hei93a, KS93a, Opp92, RWB94, TCS92]. spectrum [Sch91a]. spherical [HF93, IM92]. Spline [Wan92, Beh90, BG93, Ham94a, KB93, Liq91, Ram90, Ste94]. splines [HJ92, SV94]. splittings [Dav94]. spreading [CFT^{+}93]. sprinkler [MN91]. square [LYX^{+}93, TM92]. squared [LMV91]. squared-magnitude [LMV91]. squares [CXL94]. Stability [EN94, Lop92, Ten90, Zag91, ZL90, FLZ91, FLG92, FL94, GS94a, JNF95, KS93b, LS91b, Rua92, VB92]. Stabilizing [LSVS94, TFM93]. stable [Sim92]. Stackelberg [SMO91]. standardized [BD94]. start [VBH91]. State [Ezz94, EB91, MS93c, MK93, Pat90, Siv92]. state-variable [Pat90]. stationary [Han92, MS94a]. Statistical [Bay93, Yip93]. statistics [CH93]. steady [BH90, Siv92]. steady-state [Siv92]. Stefan [Asa92a]. stencil [Li94]. step [RTJL92, Sim92, Sim94]. stepping [Siv92]. Stepsize [And93]. stereo [Koz91]. Stochastic [AR90a, AR90b, MU90a, MU90b, Nam90a, Nam90b, Sha90a, Sha90b, Uen91, EB91, FLG92, FL94, HT90a, HT90b, Heu90a, Heu90b, Moh94, Pat91, PS94b, TT90a, TT90b, Unn90, Yak92a]. stochastic-dynamic-programming [EB91]. stock [GP94, PBL92]. stock-holding [PBL92]. stocks [Den90]. Stokes [Cha90a, VWSS94]. storage [FS94]. stratification [GG91a]. stratified [PSMM93]. stream [Raj94, RC91]. strict [Moh90]. strings [Cha93, Ebe93]. strokes [NW90]. strongly [JN91]. structural [Ben92]. Structure [LqGq93, KK91, PBL92, Sch91a, Wan94]. Structured [YS94]. studied [KK91]. studies [Kub94b, SKCW94]. study [Koč92, Kor92, LL94]. Sturm [WJ92]. Subbotin [Beh90]. Subject [Ano91e, Ano92f, Ano93e, Ano94h]. suboptimal [TFM93]. subspaces [KCS90, Sen90]. substrates [Zag92]. substructuring [Vas91]. sufficient [FLZ91]. suitability [GG91a]. sum [Mit92]. sum-of-product [Mit92]. summary [SA94]. Summation [Mur91a]. Summation-over-paths [Mur91a]. Sums [MMP90, PY94]. supersonic [CJF93]. support [KPP93, Mal93]. Surface [KM94, Knu93, BTW94, MS94b]. survival [MU90a, MU90b]. sustained [KS92b]. Sweeping [Kwo93]. Symbolic [GG91b]. symmetric [Hua91, LSVS94, Sha94, Udw92b]. system [Ge90b, Gun93, KKU91, Liq91, Mal93, Mor94, Mur91b, PL93, Sag91, Sav93, Son93, TGK91, VBH91, VB92]. Systems [And93, Ayo93, CUK92, EL91, FT94, FLZ91, FLG92, FL94, GG91b, JF91, JNF95, Koł94, Kon92, Kor92, LMS92, LSP91, LqGq93, Li94, LS91b, MM91, MS94c, MS92c, MR91, NFJ94, NA91, O’N90, Rua92, SP93, Udw91, Udw92a, UK94, VBH91, Wam92, ZXY94].

taken [BD90]. tangent [Knu93]. target [Ahm93]. Taylor [FG94a]. teacher [Cas91]. technique [Bad93, BG94c, JM93, RK93, dK94]. Techniques [Red90b, BBD91, Cic93a, Hei93b, Jac93, MM91, Vol92]. term [Kin94, Kum92a]. test [GG91a]. testing [Bha94, TA93]. tests [Cha91, ZW92]. th [FLG92, SV92b]. their [KCS90]. theorem [Arg94, FG94a, Gun94, KP91, Opp92]. theorems [AS90, ES90, GL92, LG94]. theoretic [Cha92b]. Theoretical [Bog91]. theory [ADM91, Cas94, Chu92, CH93, Eva90a, Eva90b, EHE93, ER94b, ER94a, Heu90a, Heu90b, KN90, MS93c, MK93, Opp92, PY94, Unn90]. theraphy [Rap92]. Thermally [Mei94]. Third [KB93, EN94, OT92, OT93]. Third-order [KB93, EN94, OT92, OT93]. Three [RWB94, CS92, JM93, Lop92, Mur91b, WM91]. three-dimensional [CS92, Mur91b, WM91]. three-point [Lop92]. time [BD90, BTW94, CW91, Da93c, Da93d, FL94, KPP93, Kum92b, Lin94, MS94c, NG94, Pap94, Pat91, PS94b, PY94, RTJL92, TLG90, Udw91, Udw92a, Wan93, Wu94]. time-delay [FL94]. time-dependent [MS94c, PS94b]. time-periodic [CW91]. time-reverse [NG94]. tip [Udw92a]. Toeplitz [LSVS94, Mor94, Sha94]. tokamak [VCM91]. tomographic [Kha94]. tomography [FEGL90]. tool [MU90a, MU90b]. topological [Sch91a]. topology [SW92]. toroidal [VCM91]. Torsional [enNAa94]. Tours [War91]. toxic [Pat90]. tracking [PS94b]. tradeoffs [PL93]. training [AH94]. transfer [HM94]. transform [Kha92]. transformation [Cic92, Faz90, Faz91]. transformations [DGV94, Ge90a, Kop90]. transforms [GM93]. transient [IBS94]. transmission [LL94]. transport [BMS92, PA94, Pat90]. travel [Sla93a]. traveling [Dro90, War91, War94]. treatment [LS93a, Rap92]. tree [Moh94, Pes91]. Trench [LSVS94]. triangle [Fun91]. triangularization [SH92]. triangulations [Wan92]. tubular [Faz91]. turbomachinery [JM93]. turbulence [SKCW94]. turbulent [Raj92, Raj94]. Turning [LW94, Lop92]. Two [FV91, BH93, Bar91, Da93a, Faz90, GM94, GL94, HV90, KS93b, LMS92, Mit90, ML90, MFF92, NSV92, SV94, Sim92, TM92, Zag92]. two-dimensional [BH93, Bar91, GL94, HV90, NSV92]. two-point [Faz90, GM94, KS93b, LMS92, ML90, MFF92, SV94, TM92]. two-step [Sim92]. type [Arg94, Cic92, Cic93a, Cic93b, EN94, EHE93, Gup90, LMV91, MS94b, MEZI93, MS94c, TH94, Wan92, ZL90, ZW92]. type-2 [Wan92]. types [CB94a, CB94b].

unbounded [VB92]. uncertainty [KPP93, Mal93, Sar93]. Unconstrained [Moh90]. unequally [Koh92]. unexpected [Pat92]. unified [CS93, Jol92]. uniform [Ada93b]. uniqueness [GS94a, Koz91]. university [CM90]. Unprogrammability [Gun93]. unsteady [CJF93, Ezz94, Raj92, Raj94]. until [PY94]. upper [OT92, OT93]. urine [WM91]. use [Ham94a, PZ94]. used [BG94c]. users [LL94]. Using [Hoj94, SB92, ZP92, Bar91, Bay93, BTW94, BP92, CWL92, CDW91, Eva90b, EHE93, GG91b, Hei93b, MM91, Moh90, MS92c, RBR94, RK93, Sav93, Sel94, Udw91, Udw92a, VBH91, Vol92, WN91, ZW90b]. utility [AaHM93, Mal91]. validation [RW93]. value [Bai91, CN90, Cas94, EL90, Faz90, GM94, Gup90, GK94, HSZ94, JNM92, KB93, KS93b, LMS92, LM94, ML90, MFF92, OT92, OT93, PA94, SV94, ŠNG92, SS94a, SW91, Sim92, Sim94, SV92a, SV92b, Ste94, TM92]. valued [SS91]. values [Gyö90, Lan92]. valve [VWSS94]. variable [Eva90a, EHE93, ER94b, ER94a, KB93, LqGq93, Pat90, Sim94]. variable-mesh [KB93]. variable-step [Sim94]. variables [EB91, Jol92, Mit90, SA94, VS90]. Variance [Jac93]. Variation [ZXY94, SA94, CB94a, CB94b]. Variational [CK90, LSP91, Knu91, Koč92]. varying [SS94a]. Vector [ZG90, CC94, Rug93, ZM91]. vectorized [HV90]. vehicle [Mat93]. velocity [MS94a, Raj94]. velocity-current [MS94a]. version [Mit92]. versus [HV90]. verticum [MS94c]. very [Kha90]. via [BS93, Ge91, HKLS93, RC91, RRA90, Sen90, TM92, VCM91, XYY94]. vibration [HF93]. visibility [BD94]. Voigt [LW90]. Volterra [Abd93, GL94, Mor90, Rua92, YF91, ZL90]. Volterra-series [YF91]. volume [And90]. volumes [Ano91e, Ano92f, Ano93e, Ano94h]. vs [RBR94].

water [Bog91, GGR91, MN91, PL93]. wave [enNAa94, Jac90, Reg94]. way [Mat93]. Weak [Cro93, Hua91]. weakly [Da93c, Da93d, Pap92]. weapon [Ahm93]. Weibull [BP92, Sos91]. weight [Da93b]. weighted [VS90]. weighting [CH93, Lin94]. well [ML90]. well-conditioning [ML90]. widely [BJQBQ94]. wishes [Kag91a]. within [Mur93]. Work [Tes91b, Tes91c, Kag91b]. World [Sla93b, Mus94].

Yamamoto [Arg94]. Yamamoto-type [Arg94]. zero [PY94]. zeroes [KS92a]. zeros [ZW90a].

References

Al-alwani:1993:IIM

[AaHM93] Jumah E. Al-alwani, Benjamin F. Hobbs, and Behnam B. Malakooti. An interactive integrated multiobjective optimization approach for quasiconcave / quasiconvex utility functions. Applied Mathematics and Computation, 54(2–3):241–257, March 15, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039390061I.

Azaiez:1994:PAB

[AB94] M. Naceur Azaiez and Vicki M.

Bier. Perfect aggregation in Bayesian estimation for several realistic reliablity models. Applied Mathematics and Computation, 64(2–3):139–153, September 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300394900590.

Abdalkhani:1992:NCS

[Abd92] J. Abdalkhani. The nonlinear cooling of a semi-infinite solid– implicit Runge–Kutta (IRK) methods. Applied Mathematics and Computation, 52(2–3): 233–237, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290080K.

Abdalkhani:1993:HOM

[Abd93] J. Abdalkhani. Higher order methods for solving Volterra integrodifferential equations of the first kind. Applied Mathematics and Computation, 58(1):97– 101, September 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390045G.

Argyros:1994:MMB

[AC94] Ioannis K. Argyros and Dong Chen. The midpoint method in Banach spaces and the Ptâk error estimates. Applied Mathematics and Computation, 62 (1):1–15, April 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901295.

Adams:1991:REM

[Ada91] John C. Adams. Recent enhancements in MUDPACK, a multigrid software package for elliptic partial differential equations. Applied Mathematics and Computation, 43(1):79–93, May 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391900955.

Adamatzky:1993:ICA

[Ada93a] Andrew I. Adamatzky. Implantation of cellular automata. Applied Mathematics and Computation, 55(1):49–71, April 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390005Y.

Adams:1993:MMS

[Ada93b] John C. Adams. MUDPACK-2: multigrid software for approximating elliptic partial differential equations on uniform grids with any resolution. Applied Mathematics and Computation, 53(2– 3):235–249, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390104M.

Adeff:1994:SAP

[Ade94] Sergio E. Adeff. Simulation of axisymmetric piezoelectric devices. Applied Mathematics and Computation, 65(1–3):355–369, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901880.

Ammar:1991:EIT

[ADM91] G. Ammar, W. Dayawansa, and C. Martin. Exponential interpolation: theory and numerical algorithms. Applied Mathematics and Computation, 41(3):

189–232, February 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190025I.

Adeli:1994:ACG

[AH94] H. Adeli and S. L. Hung. An adaptive conjugate gradient learning algorithm for efficient training of neural networks. Applied Mathematics and Computation, 62(1):81–102, April 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901341.

Ahmad:1993:NEC

[Ahm93] Ibrahim A. Ahmad. Nonparametric estimation of the CDF of weapon effectiveness on target performance. Applied Mathematics and Computation, 57(2– 3):289–303, September 1, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393901536.

Allen:1990:MAS

[ALMS90] Linda Allen, Truman Lewis, Clyde F. Martin, and Mark Stamp. A mathematical analysis and simulation of a localized measles epidemic. Applied Mathematics and Computation, 39(1): 61–77, September 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090122J.

Anderson:1990:GCV

[And90] Dale A. Anderson. Grid cell volume control with an adaptive grid generator. Applied Mathematics and Computation, 35(3): 209–217, February 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300390900422.

Andrus:1993:RKM

[And93] Jan Frederick Andrus. A Runge– Kutta method with stepsize control for separated systems of firstorder ODEs. Applied Mathematics and Computation, 59(2–3): 193–214, December 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390089W.

Anonymous:1990:EBa

[Ano90a] Anonymous. Editorial Board.

Applied Mathematics and Computation, 35(1):ii–iii, January 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039090088K.

Anonymous:1990:EBb

[Ano90b] Anonymous. Editorial board. Applied Mathematics and Computation, 36(1):ii–iii, March 1990.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090070J.

Anonymous:1990:EBc

[Ano90c] Anonymous. Editorial Board. Applied Mathematics and Computation, 37(1):ii–iv, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090105C.

Anonymous:1990:EBd

[Ano90d] Anonymous. Editorial Board. Applied Mathematics and Computation, 38(1):ii–iii, July 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090076F.

Anonymous:1990:EBe

[Ano90e] Anonymous. Editorial board. Applied Mathematics and Computation, 39(1):ii–iii, September 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090117L.

Anonymous:1990:EBf

[Ano90f] Anonymous. Editorial Board. Applied Mathematics and Computation, 40(1):ii–iii, November 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039090093I.

Anonymous:1990:I

[Ano90g] Anonymous. Index for 1990. Applied Mathematics and Computation, 40(3):263–265, December 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0096300308800070.

Anonymous:1991:EBa

[Ano91a] Anonymous. Editorial board. Applied Mathematics and Computation, 41(1):??, January 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190102S.

Anonymous:1991:EBb

[Ano91b] Anonymous. Editorial board. Applied Mathematics and Computation, 42(1):??, March 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391901109.

Anonymous:1991:EBc

[Ano91c] Anonymous. Editorial Board.

Applied Mathematics and Computation, 43(1):ii–iii, May 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190090A.

Anonymous:1991:EBd

[Ano91d] Anonymous. Editorial board. Applied Mathematics and Computation, 46(1):ii–iii, November 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300391900977.

Anonymous:1991:SIV

[Ano91e] Anonymous. Subject index to volumes 41–46, 1991. Applied Mathematics and Computation, 46(3):i–v, December 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300391901198.

Anonymous:1992:AMC

[Ano92a] Anonymous. Applied Mathematics and Computation. Applied Mathematics and Computation, 52(2–3):vi–viii, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290093G.

Anonymous:1992:EBa

[Ano92b] Anonymous. Editorial board. Applied Mathematics and Computation, 47(1):ii–iii, January 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290059A.

Anonymous:1992:EBb

[Ano92c] Anonymous. Editorial Board.

Applied Mathematics and Computation, 49(1):ii–iii, May 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900523.

Anonymous:1992:EBc

[Ano92d] Anonymous. Editorial Board.

Applied Mathematics and Computation, 51(1):??, September 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039290131J.

Anonymous:1992:EBd

[Ano92e] Anonymous. Editorial Board.

Applied Mathematics and Computation, 52(1):ii–iii, November 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039290094H.

Anonymous:1992:SIV

[Ano92f] Anonymous. Subject index to volumes 47–52, 1992. Applied Mathematics and Computation, 52(2–3):i–v, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290092F.

Anonymous:1993:EBa

[Ano93a] Anonymous. Editorial Board. Applied Mathematics and Computation, 53(1):ii–iii, January 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039390111Q.

Anonymous:1993:EBb

[Ano93b] Anonymous. Editorial board. Applied Mathematics and Computation, 55(1):ii–iii, April 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390001U.

Anonymous:1993:EBc

[Ano93c] Anonymous. Editorial Board. Applied Mathematics and Computation, 56(1):ii–iii, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390074O.

Anonymous:1993:EBd

[Ano93d] Anonymous. Editorial board. Applied Mathematics and Computation, 57(1):ii–iii, August 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300393900094.

Anonymous:1993:SIV

[Ano93e] Anonymous. Subject index to volumes 53–59, 1993. Applied Mathematics and Computation, 59(2–3):299–303, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390094U.

Anonymous:1994:EBa

[Ano94a] Anonymous. Editorial Board.

Applied Mathematics and Computation, 60(1):iv, January 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394902003.

Anonymous:1994:EBb

[Ano94b] Anonymous. Editorial Board.

Applied Mathematics and Computation, 61(1):ii–iii, March 1, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300394901414.

Anonymous:1994:EBc

[Ano94c] Anonymous. Editorial Board.

Applied Mathematics and Computation, 62(1):ii–iii, April 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901287.

Anonymous:1994:EBd

[Ano94d] Anonymous. Editorial Board.

Applied Mathematics and Computation, 63(1):ii–iii, June 1994.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901473.

Anonymous:1994:EBe

[Ano94e] Anonymous. Editorial board. Applied Mathematics and Computation, 64(1):ii–iii, August 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039490135X.

Anonymous:1994:EBf

[Ano94f] Anonymous. Editorial Board. Applied Mathematics and Computation, 65(1–3):ii–iii, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0096300394901600.

Anonymous:1994:EBg

[Ano94g] Anonymous. Editorial board. Applied Mathematics and Computation, 66(1):ii–iii, November 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901538.

Anonymous:1994:SIV

[Ano94h] Anonymous. Subject index to volumes 60–66, 1994. Applied Mathematics and Computation, 66(2–3):343–348, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901279.

Artzrouni:1990:SDEa

[AR90a] Marc Artzrouni and James Reneke. Stochastic Differential equations in mathematical demography: a review. Applied Mathematics and Computation, 38(1):7–21, July 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090078H.

Artzrouni:1990:SDEb

[AR90b] Marc Artzrouni and James Reneke. Stochastic differential equations in mathematical demography: a review. Applied Mathematics and Computation, 39(3):139–153, October 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090006O.

Adomian:1993:NAM

[AR93] G. Adomian and R. Rach. A new algorithm for matching boundary conditions in decomposition solutions. Applied Mathematics and Computation, 57 (1):61–68, August 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393900124.

Argyros:1990:SMG

[Arg90] Ioannis K. Argyros. The secant method in generalized Banach spaces. Applied Mathematics and Computation, 39(2): 111–121, September 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090026Y.

Argyros:1994:CTN

[Arg94] Ioannis K. Argyros. A convergence theorem for Newtonlike methods under generalized Chen–Yamamoto-type assumptions. Applied Mathematics and Computation, 61(1): 25–37, March 1, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901449.

Adomian:1994:SGN

[ARM94] G. Adomian, R. Rach, and R. E. Meyers. Solution of generic nonlinear oscillators. Applied Mathematics and Computation, 64(2– 3):167–170, September 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394900612.

Argyros:1990:CTA

[AS90] Ioannis K. Argyros and Ferenc Szidarovszky. Comparison theorems for algorithmic models. Applied Mathematics and Computation, 40(2):179–185, November 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039090131L.

Argyros:1992:MCA

[AS92] Ioannis K. Argyros and Ferenc Szidarovszky. On the monotone convergence of algorithmic models. Applied Mathematics and Computation, 48(2– 3):167–176, April 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290036Z.

Abu-Sitta:1994:NCB

[AS94a] A. M. M. Abu-Sitta. A note on a certain boundary-layer equation. Applied Mathematics and Computation, 64(1):73–77, August 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300394901406.

Abu-Sitta:1994:SBL

[AS94b] A. M. M. Abu-Sitta. On solutions of a boundary layer equation. Applied Mathematics and Computation, 63(1):95–100, June 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039490152X.

Asaithambi:1992:GMS

[Asa92a] N. S. Asaithambi. A Galerkin method for Stefan problems. Applied Mathematics and Computation, 52(2–3):239–250, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039290081B.

Asaithambi:1992:NMI

[Asa92b] N. S. Asaithambi. A numerical method for an inhomogeneous nonlinear diffusion problem. Applied Mathematics and Computation, 50(2–3):307–319, March 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039290130S.

Anuradha:1994:SBC

[ASZ94] V. Anuradha, R. Shivaji, and J. Zhu. On s-shaped bifurcation curves for multiparameter positone problems. Applied Mathematics and Computation, 65(1– 3):171–182, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300394901740.

Avgerinos:1994:OCP

[Avg94] Evgenios P. Avgerinos. Optimal control problems involving optimal economic growth with infinite horizon. Applied Mathematics and Computation, 66(2–3):

293–311, December 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901244.

Ayoubi:1993:RGF

[Ayo93] I. S. Ayoubi. The Riemann– Green function and the invariant imbedding equations for hyperbolic systems of first-order. Applied Mathematics and Computation, 55(2–3):101–129, May 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393900179.

Badiru:1993:NCS

[Bad93] Adedeji B. Badiru. A new computational search technique for AI based on the Cantor set. Applied Mathematics and Computation, 57(2–3):255–274, September 1, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300393901514.

Bai:1991:PCM

[Bai91] Yin Bai. A perturbed collocation method for boundary-value problems in differential-algebraic equations. Applied Mathematics and Computation, 45(3): 269–291, October 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/009630039190084Z.

Balakrishnan:1994:SCP

[Bal94] A. V. Balakrishnan. Shape control of plates with piezo actuators and collocated position / rate sensors. Applied Mathematics and Computation, 63(2– 3):213–234, July 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901961.

Barthelson:1991:ATD

[Bar91] Stephen H. Barthelson. Acceleration of a two-dimensional combustion code using MOSI to find the pressure and pressurecorrection fields for a SIMPLER based algorithm. Applied Mathematics and Computation, 46(3): 211–220, December 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391901176.

Bayard:1993:SPS

[Bay93] D. S. Bayard. Statistical plant set estimation using Schroederphased multisinusoidal input design. Applied Mathematics and Computation, 58(2–3): 169–198, October 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393901352.

Bardossy:1991:FSP

[BBD91] A. Bardossy, I. Bogardi, and L. Duckstein. Fuzzy set and probabilistic techniques for healthrisk analysis. Applied Mathematics and Computation, 45(3): 241–268, October 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190083Y.

Bunch:1991:DFA

[BCG91] K. J. Bunch, W. N. Cain, and R. W. Grow. A data fitting approach to series convergence acceleration. Applied Mathematics and Computation, 42(2): 189–195, March 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190050W.

Bainov:1990:PDC

[BD90] D. D. Bainov and A. B. Dishliev. Population dynamics control in regard to minimizing the time necessary for the regeneration of a biomass taken away from the population. Applied Mathematics and Computation, 39(1): 37–48, September 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090120R.

Boreux:1994:FAD

[BD94] Jean-Jacques Boreux and Lucien Duckstein. A fuzzy approach to the definition of standardized visibility in fog. Applied Mathematics and Computation, 61(2–3): 287–299, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394900515.

Behforooz:1990:MSS

[Beh90] G. Behforooz. Manifold Subbotin spline. Applied Mathematics and Computation, 40(3): 225–231, December 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300308800045.

Behforooz:1992:KPI

[Beh92]  G. Behforooz.  The not-a-

knot piecewise interpolatory cubic polynomial. Applied Mathematics and Computation, 52(1): 29–35, November 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290096J.

Benaroya:1992:REA

[Ben92] Haym Benaroya. Random eigenvalues, algebraic methods and structural dynamic models. Applied Mathematics and Computation, 52(1):37–66, November 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039290097K.

Bunch:1993:CCS

[BG93] K. J. Bunch and R. W. Grow. A compact cubic spline algorithm for converging to function minimums. Applied Mathematics and Computation, 56(1):29–34, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039390076Q.

Bomze:1994:DSE

[BG94a] Immanuel M. Bomze and Walter Gutjahr. The dynamics of self-evaluation. Applied Mathematics and Computation, 64 (1):47–63, August 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901384.

Bunch:1994:DEM

[BG94b] K. J. Bunch and R. W. Grow.

A difference equation model of partial series. Applied Mathematics and Computation, 61(2– 3):301–305, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900523.

Bunch:1994:ETI

[BG94c] K. J. Bunch and R. W. Grow. An extrapolation technique to iterate to the smallest and largest eigenvalues of an infinite-dimensional normal matrix used in function fitting. Applied Mathematics and Computation, 62(1):17–28, April 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300394901309.

Barron:1990:SMI

[BH90] R. M. Barron and M. H. Hamdan. The steady motion of an incompressible dusty gas in porous media. Applied Mathematics and Computation, 37 (3):149–167, June 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300390900536.

Barron:1993:DLB

[BH93] R. M. Barron and M. H. Hamdam. On the Darcy–Lapwood– Brinkman–Saffman dusty fluid flow models through porous media part II: applications to flow into a two-dimensional sink. Applied Mathematics and Computation, 54(1):81–97, March 1, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390034C.

Bhatti:1994:AHE

[Bha94] M. Ishaq Bhatti. Assessment of hepatic extraction of insulin by testing the small sample mean correlation coefficient. Applied Mathematics and Computation, 66(1):77–102, November 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901597.

Bo:1994:WCM

[BJQBQ94] Han Bo, Shao Ji-Qun, and Guo Bao-Qi. A widely convergent method for determining the distributed parameters of an elliptical equation. Applied Mathematics and Computation, 60(2– 3):139–146, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901007.

Bainov:1994:FOI

[BKM94] Drumi Bainov, Zdzislaw Kamont, and Emil Minchev. On first order impulsive partial differential inequalities. Applied Mathematics and Computation, 61(2–3): 207–230, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900485.

Breinbauer:1991:KMI

[BL91] M. Breinbauer and P. Lory.

The kidney model as an inverse problem. Applied Mathematics and Computation, 44(3): 195–223, August 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190058U.

Budd:1992:MTA

[BMS92] C. Budd, S. McKee, and D. Swailes. Modeling H^{+} and K^{+} transport across cell membranes. Applied Mathematics and Computation, 50(1):33–44, July 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290009P.

Bogobowicz:1991:TAM

[Bog91] Agnieszka Bogobowicz. Theoretical aspects of modeling and control of water quality in a river section. Applied Mathematics and Computation, 41 (1):35–60, January 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190105V.

Bradley:1992:GSD

[BP92] Charles E. Bradley and Tom Price. Graduating sample data using generalized Weibull functions. Applied Mathematics and Computation, 50(2–3): 115–144, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290121G.

Brezinski:1991:IJ

[Bre91] Claude Brezinski. Implementing the jackknife. Applied Mathematics and Computation, 42 (2):111–119, March 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190047Q.

Boglaev:1993:CMS

[BS93] I. P. Boglaev and V. V. Sirotkin.

Computational method for a singular perturbation problem via domain decomposition and its parallel implementation. Applied Mathematics and Computation, 56(1):71–95, June 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390080X.

Beddhu:1994:TAC

[BTW94] Murali Beddhu, Lafayette K.

Taylor, and David L. Whitfield. A time accurate calculation procedure for flows with a free surface using a modified artificial compressibility formulation. Applied Mathematics and Computation, 65(1–3):33–48, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0096300394901643.

Cabada:1994:MMF

[Cab94] Alberto Cabada. The monotone method for first-order problems with linear and nonlinear boundary conditions. Applied Mathematics and Computation, 63(2–3):163–186, July 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901937.

See erratum [Cab95].

Cabada:1995:ERM

[Cab95] Alberto Cabada. Erratum: “Remarks on the monotone method for first-order problems with linear and nonlinear boundary conditions”, Applied Math. Comput. 63(2–3), 163–186, July 1994. Applied Mathematics and Computation, 69(2–3):353–354, May 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039400147V.

See [Cab94].

Cai:1993:NEP

[Cai93]  Zhiqiang Cai.  Norm esti-

mates of product operators with application to domain decomposition. Applied Mathematics and Computation, 53(2–3): 251–276, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/009630039390105N.

Chen:1993:NHM

[CAQ93] Dong Chen, I. K. Argyros, and Q. S. Qian. A note on the Halley method in Banach spaces. Applied Mathematics and Computation, 58(2–3):215–224, October 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300393901374.

Carmody:1990:MPF

[Car90] Kevin Carmody. Modular parts of a function. Applied Mathematics and Computation, 36 (1):63–74, March 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090075E.

Casti:1991:BKT

[Cas91] John L. Casti. Bob Kalaba:

teacher, colleague, and friend. Applied Mathematics and Computation, 45(2):85–88, September 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300391900704.

Castro:1994:BTR

[Cas94]  Alfonso Castro.  Bifurcation

theory and radial solutions to elliptic boundary value problems. Applied Mathematics and Computation, 65(1–3):223–230, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901783.

Cicci:1994:SEKa

[CB94a] David A. Cicci and Gary H. Ballard. Sensitivity of an extended Kalman filter. 1. Variation in the number of observers and types of observations. Applied Mathematics and Computation, 66(2– 3):233–246, December 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901198.

Cicci:1994:SEKb

[CB94b] David A. Cicci and Gary H. Ballard. Sensitivity of an extended Kalman filter. 2. Variation in the observation error levels, observation rates, and types of observations. Applied Mathematics and Computation, 66(2–3): 247–259, December 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901201.

Cavazos-Cadena:1994:CAC

[CC94] Rolando Cavazos-Cadena. Computing the asymptotic covariance matrix of a vector of sample autocorrelations for ARMA processes. Applied Mathematics and Computation, 64(2–3):

121–137, September 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900582.

Chin:1991:FAP

[CDW91] Carolyn Chin, Lucien Duckstein, and Melanie L. Wymore. Factory automation project selection using multicriterion Qanalysis. Applied Mathematics and Computation, 46(2):107– 126, November 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190019J.

Comincioli:1993:MMP

[CFT^{+}93] V. Comincioli, D. Funaro, A. Torelli, E. D’Angelo, and P. Rossi. A mathematical model of potential spreading along neuron dendrites of cerebellar granule cells. Applied Mathematics and Computation, 59(1): 73–87, November 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390035D.

Cicci:1993:OWP

[CH93] David A. Cicci and Roger L.

Hall. Optimal weighting of a priori statistics in linear estimation theory. Applied Mathematics and Computation, 57(2–3):221– 231, September 1, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300393901488.

Chang:1990:MFE

[Cha90a] Ching Lung Chang. A mixed finite element method for the Stokes problem: an accelerationpressure formulation. Applied Mathematics and Computation, 36(2):135–146, March 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090016V.

Chang:1990:SME

[Cha90b] Feng Cheng Chang. Solution of Maxwell’s equations by matrix formulation. Applied Mathematics and Computation, 40(3): 239–262, December 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300308800069.

Chankong:1991:OTP

[Cha91] Vira Chankong. Optimality tests for partitioning and sectional search algorithms. Applied Mathematics and Computation, 44 (2):157–182, July 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190016G.

Chaitin:1992:LPSc

[Cha92a] C. J. Chaitin. LISP program-size complexity III. Applied Mathematics and Computation, 52(1): 127–139, November 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392901016.

Chaitin:1992:ITI

[Cha92b] G. J. Chaitin. Informationtheoretic incompleteness. Applied Mathematics and Computation, 52(1):83–101, November 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290099M.

Chaitin:1992:LPSa

[Cha92c] G. J. Chaitin. LISP programsize complexity. Applied Mathematics and Computation, 49 (1):79–93, May 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900578.

Chaitin:1992:LPSb

[Cha92d] G. J. Chaitin. LISP program-size complexity II. Applied Mathematics and Computation, 52(1): 103–126, November 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290100F.

Chaitin:1992:LPSd

[Cha92e] G. J. Chaitin. LISP program-size complexity IV. Applied Mathematics and Computation, 52(1): 141–147, November 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300392901027.

Chaitin:1993:NNB

[Cha93] G. J. Chaitin. A note on the number of N-bit strings with maximum complexity. Applied Mathematics and Computation, 59(1): 97–100, November 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390037F.

Chukwu:1992:MCT

[Chu92] E. N. Chukwu. Mathematical controllability theory of capital growth of nations. Applied Mathematics and Computation, 52(2– 3):317–344, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290086G.

Chukwu:1994:OCG

[Chu94] E. N. Chukwu. Optimal control of the growth of income of nations. Applied Mathematics and Computation, 62(2– 3):279–309, May 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900884.

Cicci:1992:AHT

[Cic92] David A. Cicci. Applications of the Householder transformation to Ridge-type estimation methods. Applied Mathematics and Computation, 51(2–3): 159–165, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900718.

Cicci:1993:CBP

[Cic93a] David A. Cicci. Comparison of biasing parameter computational techniques in ridge-type estimation. Applied Mathematics and Computation, 53(2–3): 111–120, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390096W.

Cicci:1993:SRT

[Cic93b] David A. Cicci. Sensitivity of ridge-type estimation methods to observation accuracy and sampling rate. Applied Mathematics and Computation, 57(2–3):233– 244, September 1, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393901499.

Ciucci:1993:NAU

[CJF93] A. Ciucci, Rhonald M. Jenkins, and Winfred A. Foster, Jr. Numerical analysis of an unsteady confined supersonic jet. Applied Mathematics and Computation, 57(2–3):245–254, September 1, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039390150D.

Calude:1994:IE

[CJZ94] C. Calude, H. Jürgensen, and M. Zimand. Is independence an exception? Applied Mathematics and Computation, 66(1): 63–76, November 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901589.

Choe:1990:VAP

[CK90] W. H. Choe and K. Kim. Variational analysis of a plasma-arc armature inside a railgun solidhydrogen-pellet accelerator for fueling fusion reactors. Applied Mathematics and Computation, 37(1):37–56, May 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090109G.

Conrad:1990:BCP

[CM90] Michael Conrad and Koichiro Matsuno. The boundary condition paradox: a limit to the university of differential equations. Applied Mathematics and Computation, 37(1):67–74, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090111F.

Cabada:1990:ESS

[CN90] A. Cabada and Juan J. Nieto.

Extremal solutions of second order nonlinear periodic boundary value problems. Applied Mathematics and Computation, 40(2): 135–145, November 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090128P.

Conrad:1993:RLR

[Con93a] Deborah Conrad. Rules, laws, and reductionism. Applied Mathematics and Computation, 56 (2–3):341–358, July 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390127Z.

Conrad:1993:B

[Con93b] Michael Conrad. Biophysicalism.

Applied Mathematics and Computation, 56(2–3):103–109, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039390117W.

Conrad:1993:FMF

[Con93c] Michael Conrad. The Fluctuon Model of force, life, and computation: a constructive analysis. Applied Mathematics and Computation, 56(2–3):203–259, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390123V.

Crombez:1993:WNC

[Cro93] G. Crombez. Weak and norm convergence of a parallel projection method in Hilbert spaces. Applied Mathematics and Computation, 56(1):35–48, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390077R.

Chan:1992:ETD

[CS92] William M. Chan and Joseph L. Steger. Enhancements of a threedimensional hyperbolic grid generation scheme. Applied Mathematics and Computation, 51(2– 3):181–205, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290073A.

Chatterjea:1993:UPC

[CS93] S. K. Chatterjea and H. M. Srivastava. A unified presentation of certain operational formulas for the Jacobi and related polynomials. Applied Mathematics and Computation, 58(1):77– 95, September 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390044F.

Chukwu:1992:GCN

[CUK92] E. N. Chukwu, Robert G. Underwood, and Louis D. Kovari. Global constrained nullcontrollability of nonlinear neutral systems. Applied Mathematics and Computation, 49 (1):95–110, May 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900589.

Chan:1991:CMT

[CW91] C. Y. Chan and Benedict M.

Wong. Computational methods for time-periodic solutions of singular semilinear parabolic problems. Applied Mathematics and Computation, 42(3): 287–312, April 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391900058.

Chan:1992:CEA

[CW92] C. Y. Chan and Benedict M.

Wong. Computational error analysis of periodic solutions for singular semilinear parabolic problems. Applied Mathematics and Computation, 50(2–3): 279–306, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290129O.

Chang:1992:NIF

[CWL92] Qianshun Chang, Yau Shu Wong, and Zhengfeng Li. New interpolation formulas of using geometric assumptions in the algebraic multigrid method. Applied Mathematics and Computation, 50(2– 3):223–254, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290127M.

Chang:1994:AEC

[CXL94] Ching L. Chang, Xiaoqing Xiang, and Jianshi Li. An analysis of the Eddy current problem by the least-squares finite element in 2-D. Applied Mathematics and Computation, 60(2– 3):179–191, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039490104X.

Chronopoulos:1992:IMN

[CZ92] A. T. Chronopoulos and Z. Zlatev. Iterative methods for nonlinear operator equations. Applied Mathematics and Computation, 51(2–3):167–180, October 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900729.

Da:1993:FEM

[Da93a] Xu Da. Finite element methods of the two nonlinear integrodifferential equations. Applied Mathematics and Computation, 58(2–3):241–273, October 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393901396.

Da:1993:NSI

[Da93b] Xu Da. Non-smooth initial data error estimates with the weight norms for the linear finite element method of parabolic partial differential equations. Applied Mathematics and Computation, 54(1):1–24, March 1, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393900319.

Da:1993:DTPa

[Da93c] Xu Da. On the discretization in time for a parabolic integrodifferential equation with a weakly singular kernel I: smooth initial data. Applied Mathematics and Computation, 57(1):1–27, August 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039390010C.

Da:1993:DTPb

[Da93d] Xu Da. On the discretization in time for a parabolic integrodifferential equation with a weakly singular kernel II: nonsmooth initial data. Applied Mathematics and Computation, 57(1):29–60, August 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300393900113.

Davis:1994:FDS

[Dav94] Stephen F. Davis. Flux difference splittings and limiters for the resolution of contact discontinuities. Applied Mathematics and Computation, 65(1–3):3–18, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901627.

Ding:1993:HOA

[DDL93] J. Ding, Q. Du, and T. Y. Li. High order approximation of the Frobenius–Perron operator. Applied Mathematics and Computation, 53(2–3):151–171, February 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039390099Z.

Dendrinos:1990:SDS

[Den90] Dimitrios S. Dendrinos. Spatially distributed stocks, deterministic evolution, and ecological balance. Applied Mathematics and Computation, 36(2):113–133, March 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039090015U.

deFalco:1994:RRR

[dFFG94] D. de Falco, M. Frontini, and L. Gotusso. Recent results on the regularization of Fourier polynomials. Applied Mathematics and Computation, 66 (1):1–8, November 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901546.

Doedel:1994:LCO

[DFM94] Eusebius J. Doedel, Mark J.

Friedman, and Anand C. Monteiro. On locating connecting orbits. Applied Mathematics and Computation, 65(1–3):231–239, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901791.

Duckstein:1993:PSI

[DGP93] Lucien Duckstein, Ambrose Goicoechea, and Eric Parent. Preface to the special issue on multicriterion decision making with engineering applications. Applied Mathematics and Computation, 54(2–3):99, March 15, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039390054I.

Daemen:1994:ISI

[DGV94] Joan Daemen, René Govaerts, and Joos Vandewalle. Invertible shift-invariant transformations on binary arrays. Applied Mathematics and Computation, 62(2–3):259–277, May 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394900876.

De:1990:AHP

[DHK90] Mitali De, Keith W. Hipel, and D. Marc Kilgour. Algorithms for hierarchical power. Applied Mathematics and Computation, 39(1):21–36, September 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090119N.

deKlerk:1994:SCA

[dK94] J. H. de Klerk. The sensitivity of a computational L_{1}- approximation technique for integral equations. Applied Mathematics and Computation, 62 (2–3):171–182, May 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394900809.

Dostal:1990:PPD

[Dos90] Zdenĕk Dostál. Projector preconditioning and domain decomposition methods. Applied Mathematics and Computation, 37(2):75–81, May 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300390900363.

Dror:1990:CAT

[Dro90] Moshe Dror. Cost allocation: the traveling salesman, binpacking, and the knapsack. Applied Mathematics and Computation, 35(2): 191–207, January 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090116K.

El-Attar:1994:CAP

[EA94] M. El-Attar. Convergence analysis for a parallel Jacobi algorithm. Applied Mathematics and Computation, 62(2–3):191–201, May 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300394900825.

Estalrich:1991:ASS

[EB91] Juan Estalrich and Nathan Buras. Alternative specifications of state variables in stochasticdynamic-programming models of reservoir operation. Applied Mathematics and Computation, 44(2):143–155, July 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190015F.

Ebeling:1993:COI

[Ebe93] Werner Ebeling. Chaos, order, and information in the evolution of strings. Applied Mathematics and Computation, 56 (2–3):131–151, July 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390119Y.

Evans:1993:ESS

[EHE93] James D. Evans, John H. Hubbell, and Valerie D. Evans. Exact series solution to the Epstein–Hubbell generalized elliptic type integral using complex variable residue theory. Applied Mathematics and Computation, 53(2–3):173–189, February 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390100S.

ElAttar:1992:FEA

[El 92] M. El Attar. Finite element analysis on distributed memory architectures. Applied Mathematics and Computation, 52(2–3): 309–316, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/009630039290085F.

Erbe:1990:BVP

[EL90] L. H. Erbe and Xinzhi Liu.

Boundary value problems for nonlinear impulsive integrodifferential equations. Applied Mathematics and Computation, 36 (1):31–50, March 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090073C.

Erbe:1991:MIM

[EL91] Lynn Erbe and Xinzhi Liu.

Monotone iterative methods for differential systems with finite delay. Applied Mathematics and Computation, 43(1):43–64, May 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300391900933.

Elabbasy:1994:OBS

[Ela94] E. M. Elabbasy. Oscillatory behavior of solutions of forced second order differential equations with alternating coefficients. Applied Mathematics and Computation, 60(1):1–13, January 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394902011.

El-Namoury:1992:FEM

[EN92] A. R. M. El-Namoury. On the finite element method for a singular problem. Applied Mathematics and Computation, 49(2– 3):239–249, June 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290027X.

El-Nahhas:1994:STO

[EN94] A. El-Nahhas. Stability of a third-order differential equation of retarded type. Applied Mathematics and Computation, 60(2– 3):147–152, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901015.

Abd-alla:1994:TWP

[enNAa94] Abo el-nour N. Abd-alla. Torsional wave propagation in an orthotropic magneto-elastic hollow circular cylinder. Applied Mathematics and Computation, 63(2–3):281–293, July 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901996.

Eno:1990:SAE

[Eno90] Larry Eno. Sensitivity analysis of experimental data. II. New coefficients. Applied Mathematics and Computation, 37(1):9–17, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039090107E.

Evans:1994:ACVb

[ER94a] James D. Evans and Thomas L.

Roberson. Applications of complex variable residue theory for evaluating irrational definite integrals IV. Applied Mathematics and Computation, 63 (1):35–57, June 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039490149X.

Evans:1994:ACVa

[ER94b] James D. Evans and Thomas L.

Roberson. Applications of complex variable residue theory to the evaluation of irrational definite integrals III. Applied Mathematics and Computation, 63(1):1–34, June 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901481.

Erlander:1990:GRT

[ES90] S. Erlander and T. E. Smith.

General representation theorems for efficient population behavior. Applied Mathematics and Computation, 36(3):173–217, April 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039090012R.

El-Sayed:1992:FDE

[ES92] A. M. A. El-Sayed. On the fractional differential equations. Applied Mathematics and Computation, 49(2–3):205–213, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290024U.

El-Sayed:1993:LDE

[ES93a] A. M. A. El-Sayed. Linear differential equations of fractional orders. Applied Mathematics and Computation, 55(1):1–12, April 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039390002V.

El-Sheikh:1993:ONC

[ES93b] M. M. A. El-Sheikh. Oscillation and nonoscillation criteria for second-order nonlinear differential equations II. Applied Mathematics and Computation, 59(2– 3):101–115, December 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390083Q.

El-Sheikh:1994:OBSb

[ES94] M. M. A. El-Sheikh. On oscillatory behavior of solutions of second order functional differential equations. Applied Mathematics and Computation, 62 (2–3):183–189, May 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900817.

El-Sheikh:1994:OBSa

[ESS94] M. M. A. El-Sheikh and R. A.

Sallam. On the oscillatory behavior of solutions of second order differential equations. Applied Mathematics and Computation, 61(1):1–16, March 1, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901422.

Evans:1990:ACV

[Eva90a] James D. Evans. Applications of complex variable residue theory to the evaluation of irrational definite integrals. I. Applied Mathematics and Computation, 39(2): 145–157, September 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300390900282.

Evans:1990:EFI

[Eva90b] James D. Evans. Evaluation of four irrational definite sine integrals using residue theory. Applied Mathematics and Computation, 36(3):163–172, April 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039090011Q.

Ezzat:1994:SSA

[Ezz94] Magdy A. Ezzat. State space approach to unsteady free convection flow through a porous medium. Applied Mathematics and Computation, 64(2–3): 191–205, September 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394900639.

Fazio:1990:NTM

[Faz90] R. Fazio. A noniterative transformation method applied to twopoint boundary-value problems. Applied Mathematics and Computation, 39(1):79–87, September 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039090123K.

Fazio:1991:ITM

[Faz91]  Riccardo Fazio.  The itera-

tive transformation method and length estimation for tubular flow reactors. Applied Mathematics and Computation, 42 (2):105–110, March 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190046P.

Frankel:1993:SNO

[FC93] Jay I. Frankel and S. Roy Choudhury. Some new observations on the classical logistic equation with heredity. Applied Mathematics and Computation, 58(2– 3):275–308, October 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390140A.

Fymat:1990:NPM

[FE90a] Alain L. Fymat and Jack I. Eisenman. A new phenomenon in medical imaging: effects of radiation coherence on spatial resolution. Applied Mathematics and Computation, 40(1):11–32, November 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039090095K.

Fymat:1990:OPC

[FE90b] Alain L. Fymat and Jack I. Eisenman. The optical principles of coincidence counting in nuclear medicine. Applied Mathematics and Computation, 40 (1):1–10, November 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090094J.

Fymat:1990:OCN

[FEGL90] Alain L. Fymat, Jack I. Eisenman, Moses A. Greenfield, and W. N. Paul Lee. Organ Characterization with nuclear-medicine tomography. Applied Mathematics and Computation, 35(2): 159–189, January 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090115J.

Fernandez-Gaucherand:1994:NRT

[FG94a] Emmanuel Fernández-Gaucherand. A note on the Ross–Taylor theorem. Applied Mathematics and Computation, 64(2–3):207– 212, September 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900647.

Fridrich:1994:RCO

[FG94b] Jiří Fridrich and James F. Geer. Reconstruction of chaotic orbits under finite resolution. Applied Mathematics and Computation, 66(2–3):129–159, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901139.

Feng:1994:PSA

[FL94] Zhao-Shu Feng and Yong-Qing Liu. Practical stability analysis for stochastic time-delay systems. Applied Mathematics and Computation, 63(2–3):265–279, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300394901988.

Flaam:1992:LCP

[Flå92]  Sjur Didrik Flåm.  Level-

constrained programming. Applied Mathematics and Computation, 47(2–3):259–265, February 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039290050B.

Feng:1992:CPS

[FLG92] Zhao-Shu Feng, Yong-Qing Liu, and Feng-Wei Guo. Criteria for practical stability in the p th mean of nonlinear stochastic systems. Applied Mathematics and Computation, 49(2– 3):251–260, June 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290028Y.

Fa-Long:1994:NNE

[FLYD94] Luo Fa-Long and Li Yan-Da.

Neural network for the exact RLS adaptive algorithm. Applied Mathematics and Computation, 60(2–3):103–111, February 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900973.

Feng:1991:NSC

[FLZ91] Zhao-Shu Feng, Yong-Qing Liu, and Huai-Zhong Zhao. Necessary and sufficient conditions for delay-independent stability of multidelay linear systems. Applied Mathematics and Computation, 46(1):23–32, November 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300391900999.

Fa-Long:1992:NNA

[FLZ92] Luo Fa-Long and Bao Zheng. Neural network approach to computing matrix inversion. Applied Mathematics and Computation, 47(2–3):109–120, February 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300392900408.

French:1990:CCH

[Fre90] Donald A. French. Computations on the Cahn–Hilliard model of solidification. Applied Mathematics and Computation, 40(1): 55–76, November 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090098N.

Friedman:1992:DHC

[Fri92] Charles N. Friedman. Detecting Hamiltonian cycles. Applied Mathematics and Computation, 47(2–3):137–143, February 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290042Y.

Finkelstein:1994:SCS

[FS94] Mark Finkelstein and Stephen Scheinberg. The storage complexity of some estimators of the median. Applied Mathematics and Computation, 60 (1):15–24, January 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039490202X.

Fariabi:1994:CPS

[FT94] Said Fariabi and Betty Travis.

On conjugate points for second order systems. Applied Mathematics and Computation, 60 (1):37–42, January 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394902046.

Funaro:1991:PAP

[Fun91] Daniele Funaro. Pseudospectral approximation of a PDE defined on a triangle. Applied Mathematics and Computation, 42 (2):121–138, March 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190048R.

Fei:1991:TEC

[FV91] Zhang Fei and Luis Vázquez. Two energy conserving numerical schemes for the Sine-Gordon equation. Applied Mathematics and Computation, 45(1): 17–30, September 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300391900874.

Garnaev:1990:OPN

[Gar90] Andrej Yu. Garnaev. Optimal policies in a nonlinear bioeconomic model of eutrophication. Applied Mathematics and Computation, 35(1):95–102, January 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039090092H.

Golberg:1990:CSP

[GB90] M. A. Golberg and H. Bowman. The conditioning of some projection methods for Fredholm and singular integral equations. Applied Mathematics and Computation, 40(2):165–178, November 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039090130U.

Golberg:1994:MAE

[GC94] Michael A. Golberg and C. S. Chen. On a method of Atkinson for evaluating domain integrals in the boundary element method. Applied Mathematics and Computation, 60(2–3): 125–138, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039490099X.

Ge:1990:FFT

[Ge90a] Renpu Ge. The filled function transformations for constrained global optimization. Applied Mathematics and Computation, 39(1):1–20, September 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090118M.

Ge:1990:FMM

[Ge90b] Renpu Ge. Finding more and more solutions of a system of nonlinear equations. Applied Mathematics and Computation, 36(1):15–30, March 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090072B.

Ge:1991:SLP

[Ge91] Renpu Ge. Solving linear programming problems via linear minimax problems. Applied Mathematics and Computation, 46(1):59–77, November 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039190101R.

Gil:1991:PTS

[GG91a] María Angeles Gil and Pedro Gil. A procedure to test the suitability of a factor for stratification in estimating diversity. Applied Mathematics and Computation, 43(3):221–229, June 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190007A.

Goel:1991:SCU

[GG91b] Narendra S. Goel and Mark D. Goodwin. Symbolic computation using L-systems. Applied Mathematics and Computation, 42(3):223–253, April 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300391900036.

Gatto:1991:ORE

[GGR91] Marino Gatto, Luca Ghezzi, and Sergio Rinaldi. The optimal reclamation of eutrophic water bodies. Applied Mathematics and Computation, 43 (2):105–115, May 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190027K.

Gupta:1994:SFO

[GK94] Chaitan P. Gupta and Ying C.

Kwong. L^{p}-solvability of a fourth-order boundary-value problem at resonance. Applied Mathematics and Computation, 65(1–3):151–159, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0096300394901724.

Gunji:1991:ALA

[GKn91] Yukio Gunji and Norio Kon-no.

Artificial life with autonomously emerging boundaries. Applied Mathematics and Computation, 43(3):271–298, June 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190012C.

Grace:1992:OTS

[GL92] S. R. Grace and B. S. Lalli. Oscillation theorems for second order neutral functional differential equations. Applied Mathematics and Computation, 51(2– 3):119–133, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290068C.

Guoqiang:1994:AEE

[GL94] Han Guoqiang and Zhang Liqing.

Asymptotic error expansion of two-dimensional Volterra integral equation by iterated collocation. Applied Mathematics and Computation, 61(2–3): 269–285, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900507.

Golberg:1993:NEC

[GM93] Michael A. Golberg and Michael McKenna. The numerical evaluation of a class of logarithmically singular integral transforms. Applied Mathematics and Computation, 57(2–3):117– 129, September 1, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393901422.

Ghelardoni:1994:EES

[GM94] Paolo Ghelardoni and Pietro Marzulli. Error estimates for shooting methods in two-point boundary value problems for second-order equations. Applied Mathematics and Computation, 60(2–3):237–248, February 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901074.

Golberg:1993:DSI

[Gol93] Michael A. Golberg. The discrete Sloan iterate for the generalized airfoil equation. Applied Mathematics and Computation, 57(2– 3):103–115, September 1, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390141Z.

Ghezzi:1994:QAN

[GP94] Luca Luigi Ghezzi and Lorenzo Peccati. Qualitative analysis of a nonlinear stock price model. Applied Mathematics and Computation, 63(2–3):123–130, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901902.

Ge:1990:GCF

[GQ90] Renpu Ge and Yongfeng Qin.

The globally convexized filled functions for global optimization. Applied Mathematics and Computation, 35(2):131–158, January 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039090114I.

Gross:1993:CEM

[Gro93] R. Steven Gross. Computationally efficient method for the solution of a singular integrodifferential equation. Applied Mathematics and Computation, 57(2–3):161–175, September 1, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300393901444.

Garner:1994:EUS

[GS94a] J. B. Garner and Randolph James Schilling. Existence, uniqueness, and stability for a class of diffusion problems. Applied Mathematics and Computation, 65(1– 3):161–169, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300394901732.

Graef:1994:AOB

[GS94b] John R. Graef and Paul W. Spikes. Asymptotic and oscillatory behavior of solutions of nonlinear neutral delay equations of arbitrary order. Applied Mathematics and Computation, 65(1– 3):209–221, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300394901775.

Gunji:1993:LPB

[GSK93] Yukio-Pegio Gunji, Shuji Shinohara, and Norio Konno. Learning processes based on incomplete identification and information generation. Applied Mathematics and Computation, 55 (2–3):219–253, May 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300393900238.

Goh:1991:AAS

[GT91] C. J. Goh and K. L. Teo. Alternative algorithms for solving nonlinear function and functional inequalities. Applied Mathematics and Computation, 41(2): 159–177, January 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190068X.

Gunji:1992:FLI

[Gun92] Yukio-Pegio Gunji. The form of life. I. It is possible but not necessary. Applied Mathematics and Computation, 47(2–3): 267–288, February 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900512.

Gunji:1993:FLU

[Gun93] Yukio-Pegio Gunji. Forms of life: Unprogrammability constitutes the outside of a system and its autonomy. Applied Mathematics and Computation, 58(1): 19–76, September 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390043E.

Gunji:1994:ALP

[Gun94] Yukio-Pegio Gunji. Autonomic life as the proof of incompleteness and Lawvere’s theorem of fixed point. Applied Mathematics and Computation, 61(2–3): 231–267, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900493.

Gupta:1990:FON

[Gup90] Chaitan P. Gupta. A fourthorder nonlinear boundary-value problem and an integral-type sign condition. Applied Mathematics and Computation, 35(3): 231–242, February 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300390900444.

Gyori:1990:SCE

[Gyö90] Istvan Györi. Sharp conditions for existence of nontrivial invariant cones of nonnegative initial values of difference equations. Applied Mathematics and Computation, 36(2):89–111, March 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039090014T.

Hamdan:1992:CAF

[Ham92] M. H. Hamdan. Computational aspects of flow through curved domains: the effect of no-slip. Applied Mathematics and Computation, 52(2–3): 189–206, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290077E.

Hamann:1994:CBS

[Ham94a] Bernd Hamann. Construction of B-spline approximations for use in numerical grid generation. Applied Mathematics and Computation, 65(1–3):295–314, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901848.

Hamdan:1994:SPF

[Ham94b] M. H. Hamdan. Single-phase flow through porous channels a review of flow models and channel entry conditions. Applied Mathematics and Computation, 62(2–3):203–222, May 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900833.

Han:1992:QEE

[Han92] Weimin Han. Quantitative error estimate in modeling the laminar stationary flow of a Bingham fluid. Applied Mathematics and Computation, 47 (1):15–24, January 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900615.

Hamdan:1991:ADL

[HB91] M. H. Hamdan and R. M. Barron. Analysis of the Darcy–Lapwood and the Darcy– Lapwood–Brinkman models: significance of the Laplacian. Applied Mathematics and Computation, 44(2):121–141, July 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190014E.

Hamdam:1993:DLB

[HB93] M. H. Hamdam and R. M. Barron. On the Darcy–Lapwood– Brinkman–Saffman dusty fluid flow models through porous media part I: models development. Applied Mathematics and Computation, 54(1):65–79, March 1, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039390033B.

Heinrichs:1991:SMM

[Hei91] Wilhelm Heinrichs. A 3D spectral multigrid method. Applied Mathematics and Computation, 41(2): 117–128, January 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190065U.

Heinrichs:1993:FEP

[Hei93a] Wilhelm Heinrichs. Finite element preconditioning for spectral multigrid methods. Applied Mathematics and Computation, 59(1):19–40, November 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S009630039390039H.

Heinrichs:1993:SMM

[Hei93b] Wilhelm Heinrichs. Spectral multigrid methods for domain decomposition problems using patching techniques. Applied Mathematics and Computation, 59(2–3):165–176, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390087U.

Heunis:1990:SDEa

[Heu90a] A. J. Heunis. On the stochastic differential equations of filtering theory. Applied Mathematics and Computation, 37 (3):185–218, June 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300390900558.

Heunis:1990:SDEb

[Heu90b] A. J. Heunis. On the stochastic differential equations of filtering theory. Applied Mathematics and Computation, 39 (3):3–36, October 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039090002K.

Hipel:1991:CCA

[HF91] Keith W. Hipel and Niall M. Fraser. Cooperation in conflict analysis. Applied Mathematics and Computation, 43 (2):181–206, May 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190033J.

Hwang:1993:AVS

[HF93] Daniel Y. Hwang and Winfred A. Foster, Jr. On axisymmetric vibration of shallow spherical shells with fixed circular cutout. Applied Mathematics and Computation, 57(2–3):205– 220, September 1, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300393901477.

Han:1992:CME

[HJ92] Weimin Han and Emery Jou. On the computation of minimalenergy splines: convergence analysis. Applied Mathematics and Computation, 47(1): 1–13, January 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290060E.

Hicks:1993:ANC

[HKLS93] D. L. Hicks, K. L. Kuttler, L. M.

Liebrock, and C. E. Southwell. Answering von Neumann’s conjecture on convergence of averages and analyzing the von Neumann–Richtmyer scheme via a material averaging method. Applied Mathematics and Computation, 55(2–3):265–295, May 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390025A.

Hicks:1991:PAI

[HL91] Darrell L. Hicks and Lorie M.

Liebrock. Parallel algorithms for implicit material dynamics schemes. Applied Mathematics and Computation, 46(3): 179–209, December 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391901165.

Hicks:1994:EEE

[HLW94] D. L. Hicks, L. M. Liebrock, and Y. Wen. Eigenform error estimates of computational material dynamics with shocks. 1.1. Hooke’s law materials. Applied Mathematics and Computation, 66(2–3):181–216, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901163.

Hady:1994:MCR

[HM94] F. M. Hady and R. A. Mohamed. Mixed convectionradiation interaction heat transfer in boundary-layer over a semiinfinite flat plate embedded in porous medium with a transfer magnetic field. Applied Mathematics and Computation, 66(2– 3):103–128, December 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901120.

Hojati:1994:ULP

[Hoj94] Mehran Hojati. Using linear programming to seek the optimum combination of investment. Applied Mathematics and Computation, 66(2–3):175–179, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300394901155.

Howell:1994:ECE

[How94] Gary W. Howell. Efficient computation of eigenvalues of randomly generated matrices. Applied Mathematics and Computation, 66(1):9–24, November 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901554.

Hosea:1994:EEC

[HS94] M. E. Hosea and L. F. Shampine. Estimating the error of the classic Runge–Kutta Formula. Applied Mathematics and Computation, 66(2–3):217–226, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901171.

Harris:1994:SMR

[HSZ94] G. Harris, P. Sellers, and B. Zinner. Some multiplicity results for a class of nonlinear boundary value problems. Applied Mathematics and Computation, 65(1– 3):195–208, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300394901767.

Habib:1990:ISNa

[HT90a] M. K. Habib and A. Thavaneswaran. Inference for stochastic neuronal models. Applied Mathematics and Computation, 38(1):51–73, July 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090080M.

Habib:1990:ISNb

[HT90b] M. K. Habib and A. Thavaneswaran. Inference for stochastic neuronal models. Applied Mathematics and Computation, 39(3):183–205, October 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039090008Q.

Huang:1991:CAM

[Hua91] Wei-Zhang Huang. Convergence of Algebraic multigrid methods for symmetric positive definite matrices with weak diagonal dominance. Applied Mathematics and Computation, 46(2): 145–164, November 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190022F.

Huss:1990:ALA

[Hus90] Ronald E. Huss. An ADA library for automatic evaluation of derivatives. Applied Mathematics and Computation, 35(2): 103–123, January 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090112G.

Holter:1990:CVM

[HV90] Bill Holter and George Vanden-Berghe. A comparison of vectorized methods for solving the two-dimensional diffusion equation: multigrid versus polynomial preconditioned conjugate gradient. Applied Mathematics and Computation, 40(1):77– 103, November 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/009630039090099O.

Issa:1994:AHF

[IBS94] Camille A. Issa, Krishnan Balasubramaniam, and Kanthikannan Srirengan. Adaptive hpfinite element for transient analysis. Applied Mathematics and Computation, 65(1–3):315–333, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901856.

Iakovidis:1992:IPE

[IM92] Ilias Iakovidis and Clyde F. Martin. The inverse problem in electrocardiography on a spherical model. Applied Mathematics and Computation, 48(2– 3):187–217, April 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900383.

Ioakimidis:1992:COS

[Ioa92] Nikolaos I. Ioakimidis. Computation of the orders of singularity of sectionally analytic functions. Applied Mathematics and Computation, 48(1):13–19, March 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039290015S.

Ioakimidis:1991:NEC

[IPP91] N. I. Ioakimidis, E. A. Perdios, and K. E. Papadakis. Numerical estimation of the coefficient of the homogeneous Riemann– Hilbert problem on the basis of boundary data. Applied Mathematics and Computation, 41 (1):21–33, January 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190104U.

Ismail:1993:EES

[IST93] Fudziah Bte Ismail, Mohamed B. Suleiman, and Bachok B. Taib. Existence of h ^{2}-expansion for some extrapolation methods for solving autonomous higher order ODEs. Applied Mathematics and Computation, 57 (1):69–83, August 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393900135.

Jackson:1990:EES

[Jac90] Dennis E. Jackson. Error estimates for the semidiscrete finite-element approximation of a damped wave equation with a smoothing property. Applied Mathematics and Computation, 35(3):243–261, February 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300390900455.

Jackson:1992:EES

[Jac92] D. E. Jackson. Error estimates for the semidiscrete Galerkin approximations of the FitzHugh– Nagumo equations. Applied Mathematics and Computation, 50(1):93–114, July 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290013Q.

Jacobson:1993:VBR

[Jac93] Sheldon H. Jacobson. Variance and bias reduction techniques for the harmonic gradient estimator. Applied Mathematics and Computation, 55(2–3):153–186, May 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039390019B.

Jodar:1991:IDM

[JF91] L. Jódar and M. Legua Fernandez. An implicit difference method for the numerical solution of coupled systems of partial differential equations. Applied Mathematics and Computation, 46(2):127–134, November 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190020N.

Ji:1994:ALE

[Ji94] Jun Ji. An alternative limit expression of Drazin inverse and its application. Applied Mathematics and Computation, 61(2–3): 151–156, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900442.

Jiang:1994:MFP

[Jia94] Chang Jin Jiang. A method for finding the periodic solution of autonomous ordinary differential equations. Applied Mathematics and Computation, 66(2–3): 161–173, December 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901147.

Jimenez:1994:DDC

[Jim94] Salvador Jiménez. Derivation of the discrete conservation laws for a family of finite difference schemes. Applied Mathematics and Computation, 64 (1):13–45, August 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901376.

Jing:1994:ABC

[Jin94] Chen Jing. Alternating block Crank–Nicolson method for the 3-D heat equation. Applied Mathematics and Computation, 66(1): 41–61, November 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901570.

Jingxian:1992:IMC

[JL92] Sun Jingxian and Liu Lishan. Iterative method for coupled quasisolutions of mixed monotone operator equations. Applied Mathematics and Computation, 52(2– 3):301–308, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290084E.

Jodar:1993:SSO

[JL93] Lucas Jódar and Matilde Legua.

Solving second-order matrix differential equations with regular singular points avoiding the increase of the problem dimension. Applied Mathematics and Computation, 53(2–3): 191–205, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390101J.

Jenkins:1993:ICT

[JM93] R. M. Jenkins and D. A. Moore.

An inverse calculation technique for quasi-three-dimensional turbomachinery cascades. Applied Mathematics and Computation, 57(2–3):197–204, September 1, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300393901466.

Jodar:1993:NMM

[JMV93] L. Jódar, J. L. Morera, and R. J. Villanueva. Numerical multisteps matrix methods for Y′′ = f(t, Y). Applied Mathematics and Computation, 59(2– 3):257–274, December 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390092S.

Jodar:1991:ESC

[JN91] L. Jódar and E. Navarro. Exact solution of a class of strongly coupled Riccati matrix differential equations. Applied Mathematics and Computation, 43 (2):165–173, May 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190031H.

Jodar:1995:ECS

[JNF95] L. Jódar, E. Navarro, and M. V. Ferrer. Erratum: a correction on the stability of implicit higher order difference systems. Applied Mathematics and Computation, 71(1):91–96, August 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039500060U.

See [NFJ94].

Jodar:1992:CFS

[JNM92] L. Jódar, E. Navarro, and J. L.

Morera. A closed-form solution of singular regular higher-order difference initial and boundary value problems. Applied Mathematics and Computation, 48(2– 3):153–166, April 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290035Y.

Jodar:1992:RAS

[Jód92] Lucas Jódar. Rational approximate solutions and error bounds for the nonsymmetric Riccati matrix differential equation. Applied Mathematics and Computation, 50(1):23–31, July 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290008O.

Jolayemi:1992:UAS

[Jol92] Joel K. Jolayemi. A unified approximation scheme for the convolution of independent binomial variables. Applied Mathematics and Computation, 49(2– 3):269–297, June 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900305.

Jayakumar:1993:CMS

[JR93] J. Jayakumar and N. Ramanujam. A computational method for solving singular perturbation problems. Applied Mathematics and Computation, 55 (1):31–48, April 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390004X.

Jimenez:1990:AFN

[JV90] Salvador Jiménez and Luis Vázquez. Analysis of four numerical schemes for a nonlinear Klein–Gordon equation. Applied Mathematics and Computation, 35(1):61–94, January 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090091G.

Jodar:1994:ASC

[JV94] Lucas Jódar and R. J. Villanueva. Analytic solutions of coupled second-order delay differential equations. Applied Mathematics and Computation, 62 (2–3):235–247, May 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039490085X.

Kagiwada:1991:BWB

[Kag91a] Harriet H. Kagiwada. Best wishes to Bob on his 65th birthday. Applied Mathematics and Computation, 45(2):91–92, September 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039190072U.

Kagiwada:1991:KWA

[Kag91b] Harriet H. Kagiwada. Kalaba’s work in applied mathematics: modeling, optimization, and inverse problems. Applied Mathematics and Computation, 45(2): 93–97, September 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190073V.

Kailath:1991:ROD

[Kai91] T. Kailath. Remarks on the origin of the displacement-rank concept. Applied Mathematics and Computation, 45(2):193– 206, September 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391900807.

Kadalbajoo:1993:TOV

[KB93] Mohan K. Kadalbajoo and Rajesh K. Bawa. Third-order variable-mesh cubic spline methods for singularly-perturbed boundary-value problems. Applied Mathematics and Computation, 59(2–3):117–129, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039390084R.

Kirby:1991:ELR

[KCK91] Kevin G. Kirby, Michael Conrad, and Roberto R. Kampfner. Evolutionary learning in reactiondiffusion neurons. Applied Mathematics and Computation, 41(3): 233–263, February 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190026J.

Kwak:1990:PNG

[KCS90] Do Young Kwak, U. Jin Choi, and S. Son. Prime number grid scheme to approximate function spaces with their low dimensional subspaces. Applied Mathematics and Computation, 40(3): 233–237, December 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300308800057.

Konno:1991:MCA

[KG91] Norio Konno and Yukio Gunji. Mathematical construction of an autonomous artificial life. Applied Mathematics and Computation, 46(1):33–58, November 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300391901002.

Khanh:1990:NEV

[Kha90] Bui Doan Khanh. A numerical evaluation of very oscillatory infinite integrals by asymptotics. Applied Mathematics and Computation, 40(2):117–124, November 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039090126N.

Khanh:1991:CFD

[Kha91a] Bui Doan Khanh. A computation of Fermi–Dirac integrals by asymptotics and the Hermite corrector formula. Applied Mathematics and Computation, 41 (1):61–68, January 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190106W.

Khanh:1991:NRD

[Kha91b] Bui Doan Khanh. A numerical resolution of differentialalgebraic equations. Applied Mathematics and Computation, 44(3):261–269, August 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190062R.

Khanh:1992:NEH

[Kha92] Bui Doan Khanh. A numerical evaluation of the Hankel transform of bandlimited functions. Applied Mathematics and Computation, 48(2–3):83–100, April 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039290032V.

Khanh:1994:CTR

[Kha94] Bui Doan Khanh. A computation for tomographic reconstruction from projections. Applied Mathematics and Computation, 64(2–3):79–91, September 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039490054X.

Kincanon:1991:OSC

[Kin91] Eric Kincanon. An orthogonal set composed from the functions e^{nx}, n ≥ 0, x ≤ 0. Applied Mathematics and Computation, 41(1):69–75, January 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190107X.

Kincanon:1993:ASG

[Kin93] Eric Kincanon. Approximate solution to the Gel’Fand–Levitan equation. Applied Mathematics and Computation, 53(2–3): 121–128, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390097X.

Kincanon:1994:SST

[Kin94] Eric Kincanon. A series solution for the GV ψ_{0} term of the Born series. Applied Mathematics and Computation, 66(2–3): 227–232, December 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039490118X.

Kubica:1990:ACA

[KK90] Krystian Kubica and Janina Kuczera. An application of cellular automata to model a lipid membrane. Applied Mathematics and Computation, 39(1): 49–59, September 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090121I.

Kubica:1991:EAC

[KK91] Krystian Kubica and Janina Kuczera. The effect of the alkyl chain length of modifier molecules on membrane structure studied by the cellularautomaton method. Applied Mathematics and Computation, 44(2):183–193, July 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190017H.

Kalaba:1990:LPS

[KKM90] Robert Kalaba, Moon Kim, and James E. Moore II. Linear programming and simple associative memories. Applied Mathematics and Computation, 40(3):

203–214, December 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300308800021.

Kagiwada:1991:KAM

[KKU91] Harriet H. Kagiwada, Julia K. Kagiwada, and Sueo Ueno. Kalaba’s associative memories for system identification. Applied Mathematics and Computation, 45(2):135–142, September 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190076Y.

Khamayseh:1994:SGG

[KM94] Ahmed Khamayseh and C. Wayne Mastin. Surface grid generation based on elliptic PDE models. Applied Mathematics and Computation, 65(1–3):253–264, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901813.

Krohn:1991:IML

[KMR^{+}91] I. Krohn, S. Moltzahn, J. Rosenmüller, P. Sudhölter, and H.-M. Wallmeier. Implementing the modified LH algorithm. Applied Mathematics and Computation, 45(1): 31–72, September 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0096300391900885.

Kaneko:1990:AAT

[KN90] Hideaki Kaneko and Richard Noren. An application of approximation theory to the numerical solution of a Fredholm equation with a mild singularity. Applied Mathematics and Computation, 37(1):57–66, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090110O.

Knupp:1991:DVG

[Knu91] Patrick M. Knupp. The direct variational grid generation method extended to curves. Applied Mathematics and Computation, 43(1):65–78, May 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391900944.

Knupp:1993:SGG

[Knu93] Patrick M. Knupp. Surface grid generation in the tangent plane. Applied Mathematics and Computation, 59(1):41–62, November 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S009630039390040L.

Kocvara:1992:ASL

[Koč92] Michal Kočvara. An algebraic study of a local multigrid method for variational problems. Applied Mathematics and Computation, 51(1):17–41, September 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039290133L.

Kohlbeck:1992:MNI

[Koh92] F. Kohlbeck. A method for the numerical integration of unequally spaced data. Applied Mathematics and Computation, 50(1):45–49, July 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290010X.

Kolowrocki:1994:LRF

[Koł94] Krzysztof Kołowrocki. Limit reliability functions of some series-parallel and parallel-series systems. Applied Mathematics and Computation, 62(2– 3):129–151, May 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900787.

Konno:1992:FML

[Kon92] Norio Konno. A formalism of multidimensional living systems based on nonstandard analysis. Applied Mathematics and Computation, 49(2–3):231–238, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039290026W.

Kopf:1990:IMP

[Kop90] Christoph Kopf. Invariant measures for piecewise linear transformations of the interval. Applied Mathematics and Computation, 39(2):123–144, September 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039090027Z.

Korman:1992:CAS

[Kor92] Philip Korman. A computerassisted study of periodic parabolic systems. Applied Mathematics and Computation, 50(2–3): 203–221, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290126L.

Kozera:1991:EUP

[Koz91] Ryszard Kozera. Existence and uniqueness in photometric stereo. Applied Mathematics and Computation, 44(1):1–103, July 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391900014.

Kravvaritis:1991:ETO

[KP91] Dimitrios Kravvaritis and Nikolaos S. Papageorgiou. An existence theorem for an optimal economic growth problem with infinite planning horizon. Applied Mathematics and Computation, 44(3):249–259, August 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190061Q.

Kravvaritis:1993:ESP

[KPP93] Dimitrios Kravvaritis, George Pantelides, and Nikolaos S. Papageorgiou. Existence of support prices for a discrete time growth model with uncertainty. Applied Mathematics and Computation, 53(1):27–43, January 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390114T.

Kuznetsov:1991:NAF

[KR91] Yu. A. Kuznetsov and S. Rinaldi. Numerical analysis of the flip bifurcation of maps. Applied Mathematics and Computation, 43(3):231–236, June 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190008B.

Khamayseh:1992:EBC

[KS92a] Ahmed K. Khamayseh and R. Shivaji. Evolution of bifurcation curves for semipositone problems when nonlinearities develop multiple zeroes. Applied Mathematics and Computation, 52(2–3):173–188, December 1992.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290076D.

Kriuchkov:1992:SSO

[KS92b] V. N. Kriuchkov and M. Solomonovich.

Self-sustained oscillation in economics. Applied Mathematics and Computation, 49(1):63–77, May 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300392900567.

Karageorghis:1993:CSC

[KS93a] A. Karageorghis and S. Sivaloganathan. A Chebyshev spectral collocation method for the computation of buoyancyinduced flows. Applied Mathematics and Computation, 55 (2–3):297–308, May 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390026B.

Katti:1993:SCS

[KS93b] C. P. Katti and S. Sivaloganathan. On the stability of a classical second order method for solving a class of two-point boundary value problems. Applied Mathematics and Computation, 59(2–3):247–255, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039390091R.

Kalaba:1991:SNE

[KT91] R. Kalaba and L. Tesfatsion. Solving nonlinear equations by adaptive homotopy continuation. Applied Mathematics and Computation, 41(2):99–115, January 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039190064T.

Kalaba:1992:CM

[KU92] R. Kalaba and F. Udwadia. On constrained motion. Applied Mathematics and Computation, 51(1):85–86, September 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290135N.

Kubica:1994:CAM

[Kub94a] Krystian Kubica. The cellular automata model for lipid membranes. Applied Mathematics and Computation, 63(2– 3):187–199, July 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901945.

Kubica:1994:CAS

[Kub94b] Krystian Kubica. Cellular automata studies of cholesterol effect on gramicidin ion conductivity. Applied Mathematics and Computation, 64(2–3): 101–114, September 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900566.

Kumar:1992:OER

[Kum92a] Kuldeep Kumar. Optimum exponential regression with one nonlinear term. Applied Mathematics and Computation, 50 (1):51–58, July 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290011O.

Kumar:1992:PAA

[Kum92b] Kuldeep Kumar. Padé approximation and its application in time series analysis. Applied Mathematics and Computation, 48(2– 3):139–151, April 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290034X.

Kwong:1993:SAI

[Kwo93] Man Kam Kwong. Sweeping algorithms for inverting the discrete Ginzburg–Landau operator. Applied Mathematics and Computation, 53(2–3):129–150, February 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039390098Y.

Lainiotis:1994:NDA

[LAK94] D. G. Lainiotis, N. D. Assimakis, and S. K. Katsikas. New doubling algorithm for the discrete periodic Riccati Equation. Applied Mathematics and Computation, 60(2–3):265–283, February 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901090.

Langerholc:1992:CFS

[Lan92] John Langerholc. Calculation of forward-scattering amplitudes from a focused beam: exact values for the Gaussian parameters. Applied Mathematics and Computation, 48(2–3):101–137, April 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039290033W.

Lapsa:1992:RLM

[Lap92] P. M. Lapsa. Resolution-limited measure and dimension. Applied Mathematics and Computation, 52(1):1–27, November 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290095I.

Lu:1994:DNS

[LF94] Xin Lu and Wei Feng. Dynamics and numerical simulations of food-chain populations. Applied Mathematics and Computation, 65(1–3):335–344, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901864.

Lalli:1994:OTS

[LG94] B. S. Lalli and S. R. Grace. Oscillation theorems for second order neutral difference equations. Applied Mathematics and Computation, 62(1):47–60, April 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901325.

Li:1994:NMS

[Li94] Xuefeng Li. A numerical method for systems of hyperbolic conservation laws with single stencil reconstructions. Applied Mathematics and Computation, 65(1– 3):125–140, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300394901708.

Lin:1994:FDS

[Lin94] Yanping Lin. Finite difference solutions for parabolic equations with the time weighting initial conditions. Applied Mathematics and Computation, 65(1–3): 49–61, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300394901651.

Liqing:1991:AEG

[Liq91] Zhang Liqing. Asymptotic expansion and geometric properties of the spline collocation periodic solution of an ODE system. Applied Mathematics and Computation, 42(3):209–221, April 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300391900025.

Liu:1994:BPS

[Liu94]  James H. Liu.  Bounded

and periodic solutions of differential equations in Banach space. Applied Mathematics and Computation, 65(1–3):141–150, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901716.

Lin:1994:QSH

[LL94] Xiaodong Lin and Xinzhi Liu. Qualitative study of an HIV transmission model among intravenous drug users. Applied Mathematics and Computation, 60(2– 3):215–236, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901066.

Lazer:1994:SEB

[LM94] A. C. Lazer and P. J. McKenna.

A singular elliptic boundary value problem. Applied Mathematics and Computation, 65(1–3):183– 194, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901759.

Lakshmikantham:1992:IDS

[LMS92] V. Lakshmikantham, K. N.

Murty, and S. Sivasundaram. Impulsive differential systems for two-point boundary value problems. Applied Mathematics and Computation, 50(2–3): 157–166, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290123I.

Lepschy:1991:HTI

[LMV91] A. Lepschy, G. A. Mian, and U. Viaro. Hermite-type interpolation of squared-magnitude functions. Applied Mathematics and Computation, 43(3): 265–270, June 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190011B.

Lopez:1992:STP

[Lop92] L. Lopez. Stability of a threepoint scheme for linear second order singularly perturbed BVPs with turning points. Applied Mathematics and Computation, 52(2–3):279–300, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290083D.

Li-qing:1993:SVH

[LqGq93] Zhang Li-qing and Han Guoqiang. Structure variable homotopy methods for solving nonlinear systems. Applied Mathematics and Computation, 59(2–3): 177–191, December 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390088V.

Lakshmikantham:1991:AIA

[LS91a] V. Lakshmikantham and Yong Sun. Applications of interval analysis to minimal and maximal solutions of differential equations. Applied Mathematics and Computation, 41(1):77–87, January 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039190108Y.

Liu:1991:GAS

[LS91b] Dan Liu and Ferenc Szidarovszky. Global asymptotical stability of dynamic systems with modified contractions. Applied Mathematics and Computation, 43 (3):237–240, June 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190009C.

Ladde:1993:NTR

[LS93a] G. S. Ladde and M. Sambandham. Numerical treatment of random polynomials. Applied Mathematics and Computation, 55(1):13–30, April 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390003W.

Liu:1993:NFM

[LS93b] K. Liu and R. Skelton. A new formulation of Q-Markov covariance equivalent realization. Applied Mathematics and Computation, 53(1):83–95, January 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390116V.

Lenhart:1991:VAC

[LSP91] S. Lenhart, S. Stojanovic, and V. Protopopescu. Variational approach for competitive systems with obstacles. Applied Mathematics and Computation, 42(1):65–83, March 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391901132.

Lakshmikantham:1994:CRP

[LSS94] V. Lakshmikantham, S. K. Sen, and S. Sivasundaram. Concise row-pruning algorithm to invert a matrix. Applied Mathematics and Computation, 61(1): 17–24, March 1, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901430.

Lakshmikantham:1994:STA

[LSVS94] V. Lakshmikantham, S. K. Sen, V. C. Venkaiah, and S. Sivasundaram. Stabilizing Trench’s algorithm to invert symmetric Toeplitz matrices. Applied Mathematics and Computation, 60(2– 3):249–263, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901082.

Lether:1990:ANC

[LW90] F. G. Lether and P. R. Wenston. An algorithm for the numerical computation of the Voigt function. Applied Mathematics and Computation, 35(3): 277–289, February 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300390900477.

Li:1994:HOT

[LW94] T. Y. Li and Xiaoshen Wang. Higher order turning points. Applied Mathematics and Computation, 64(2–3):155–166, September 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300394900604.

Liu:1992:MIN

[LWX92] Xinzhi Liu, Yau Shu Wong, and Ji Xingzhi. Monotone iterations for numerical solutions of nonlinear elliptic partial differential equations. Applied Mathematics and Computation, 50 (1):59–91, July 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290012P.

Lalli:1993:OHE

[LYC93] B. S. Lalli, Y. H. Yu, and B. T. Cui. Oscillation of hyperbolic equations with functional arguments. Applied Mathematics and Computation, 53(2–3): 97–110, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390095V.

Li:1993:LSF

[LYX^{+}93] J. S. Li, Z. Y. Yu, X. Q. Xiang, W. P. Ni, and C. L. Chang. Leastsquare finite element method for electromagnetic fields in 2-D. Applied Mathematics and Computation, 58(2–3):143–167, October 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039390134Z.

MacLeod:1993:NRS

[Mac93] Allan J. MacLeod. A note on the Randles–Sevcik function from electrochemistry. Applied Mathematics and Computation, 57(2– 3):305–310, September 1, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300393901547.

Madan:1990:INM

[Mad90] Ved P. Madan. Interval Newton method: Hansen–Greenberg approach — some procedural improvements. Applied Mathematics and Computation, 35(3): 263–276, February 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300390900466.

Malakooti:1991:GDM

[Mal91] Behnam B. Malakooti. A generalized decomposable multiattribute utility function. Applied Mathematics and Computation, 44(3):271–285, August 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190063S.

Malakooti:1993:DSS

[Mal93] Behnam B. Malakooti. A decision support system for discrete multiple-criteria problems: certainty, uncertainty, and hierarchical. Applied Mathematics and Computation, 54(2–3): 131–166, March 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390056K.

Matsuno:1993:BFC

[Mat93] Koichiro Matsuno. Being free from ceteris paribus: a vehicle for founding physics on biology rather than the other way around. Applied Mathematics and Computation, 56(2– 3):261–279, July 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390124W.

Malakooti:1994:SDA

[MDR94] B. Malakooti, Lucien Duckstein, and A. Ravindran. Screening discrete alternatives with imprecisely assessed additive multiattribute functions. Applied Mathematics and Computation, 60(1):69–87, January 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394902070.

Meir:1993:FEA

[Mei93] A. J. Meir. Finite element analysis of magnetohydrodynamic pipe flow. Applied Mathematics and Computation, 57(2–3):177– 196, September 1, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300393901455.

Meir:1994:TCM

[Mei94] A. J. Meir. Thermally coupled magnetohydrodynamics flow. Applied Mathematics and Computation, 65(1–3):79–94, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0096300394901678.

Mostafa:1992:MGS

[MES92] Magdi S. Mostafa and Khaled M. F. El-Sayed. Matrixgeometric solution of a multiserver queue with Markovian group arrivals and Coxian servers. Applied Mathematics and Computation, 49(2– 3):177–196, June 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290022S.

Moustafa:1994:ISQ

[MES94] Magdi S. Moustafa and Khaled M. F. El-Sayed. Iterative solution of the M^{[x]}/C_{2}/S/N queues. Applied Mathematics and Computation, 66(2–3):313–323, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300394901252.

Mohsen:1993:HAF

[MEZI93] A. Mohsen, H. El-Zoheiry, and L. Iskandar. A highly accurate finite-difference scheme for a Boussinesq-type equation. Applied Mathematics and Computation, 55(2–3):201–212, May 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393900216.

Murty:1992:STP

[MFF92] K. N. Murty, D. W. Fausett, and L. V. Fausett. Solution of two-point boundary value problems involving Kronecker products. Applied Mathematics and Computation, 50(2–3): 145–156, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290122H.

McCord:1993:MAA

[MFS93] Mark R. McCord, Oscar Franzese, and Xiao Duan Sun. Multicriteria analysis of aeromedical fleet expansion. Applied Mathematics and Computation, 54(2–3): 101–129, March 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390055J.

Martinez:1992:AMI

[MGL92] Ignacio Martínez, María Angeles Gil, and María Teresa López. Analysis of mutual information measures in cluster sampling. Applied Mathematics and Computation, 52(2–3):389–402, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039290089J.

Mitra:1990:NMM

[Mit90] A. K. Mitra. A neighborhood magnification method for solving implicit equations in two variables. Applied Mathematics and Computation, 37(2):97–110, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300390900385.

Mitra:1992:VMB

[Mit92] A. K. Mitra. A c-version of minimal Boolean sum-of-product forms. Applied Mathematics and Computation, 47(2–3):91– 108, February 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300392900394.

Mishra:1993:TTLb

[MK93] R. K. Mishra and Satish Kumar. Toward a theory of living state — Part II. Applied Mathematics and Computation, 56(2–3):167–176, July 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390121T.

Murty:1990:DWC

[ML90] K. N. Murty and P. V. S.

Lakshmi. On Dichotomy and well-conditioning in two-point boundary-value problems. Applied Mathematics and Computation, 38(3):179–199, August 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090022U.

Martin:1991:DON

[MM91] C. Martin and J. Miller. Discrete obsevability of nonlinear systems using continuation techniques. Applied Mathematics and Computation, 44(3):241–248, August 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039190060Z.

Miller:1994:SPM

[MM94] J. Miller and C. Martin. Some properties of mappings induced by Dirichlet polynomials. Applied Mathematics and Computation, 64(1):1–11, August 1994.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901368.

Martin:1990:SEP

[MMP90] C. Martin, J. Miller, and K. Pearce. Sums of exponentials with polynomial coefficients. Applied Mathematics and Computation, 40(1):33–39, November 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039090096L.

Martin:1991:AMW

[MN91] Clyde F. Martin and John Newman. Analytic models of water loss in sprinkler irrigation. Applied Mathematics and Computation, 43(1):19–41, May 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300391900922.

Mazza:1992:MMA

[MO92] Christian Mazza and Neil O’Connell. Microscopic and macroscopic aspects of epidemics. Applied Mathematics and Computation, 47(2– 3):237–258, February 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300392900497.

Mitra:1993:MPC

[MO93] Arun K. Mitra and Justin Ortiz.

A massive parallel computation in Boolean logic. Applied Mathematics and Computation, 57(2– 3):131–160, September 1, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393901433.

Mohd:1990:UGO

[Moh90] Ismail Bin Mohd. Unconstrained global optimization using strict complementary slackness. Applied Mathematics and Computation, 36(2):75–87, March 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090013S.

Mohd:1994:IEM

[Moh94] Ismail Bin Mohd. Interval elimination method for stochastic spanning tree problem. Applied Mathematics and Computation, 66(2–3):325–341, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901260.

Morhac:1990:FEF

[Mor90] Miroslav Morháč. Fast errorfree algorithm for the determination of kernels of the periodic Volterra representation. Applied Mathematics and Computation, 38(2):87–100, July 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090050D.

Morhac:1994:EFL

[Mor94] Miroslav Morháč. An error-free Levison algorithm to solve integer Toeplitz system. Applied Mathematics and Computation, 61(2–3):135–149, March 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900434.

Mosetti:1990:PIM

[Mos90] Renzo Mosetti. On a possible interpretation of market dynamics based on the internal model principle. Applied Mathematics and Computation, 37(1):1–8, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039090106D.

Mandel:1990:MFC

[MP90] Jan Mandel and Seymour V. Parter. On the multigrid F -cycle. Applied Mathematics and Computation, 37(1):19–36, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090108F.

Muratori:1991:ILS

[MR91] Simona Muratori and Sergio Rinaldi. On the invertibility of Leslie systems. Applied Mathematics and Computation, 41(2): 179–187, January 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190069Y.

Martin:1992:NEG

[MS92a] Clyde Martin and Mark Stamp.

A note on the error in Gaussian quadrature. Applied Mathematics and Computation, 47 (1):25–35, January 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900626.

Morgan:1992:CPP

[MS92b] Alexander P. Morgan and Andrew J. Sommese. Coefficientparameter polynomial continuation. Applied Mathematics and Computation, 51(2– 3):207, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290074B.

Morgan:1992:CAS

[MS92c] Alexander P. Morgan and Andrew J. Sommese. Computing all solutions to polynomial systems using homotopy continuation.

Applied Mathematics and Computation, 51(2–3):209, October 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039290075C.

Martin:1993:NEF

[MS93a] Clyde Martin and Victor Shubov. Natural exponential families of probability distributions and exponential-polynomial approximation. Applied Mathematics and Computation, 59(2–3): 275–297, December 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390093T.

Martin:1993:PMA

[MS93b] Clyde Martin and Victor Shubov. Probability measures, Appel polynomials and polynomial approximation. Applied Mathematics and Computation, 53(2–3): 277–298, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390106O.

Mishra:1993:TTLa

[MS93c] R. K. Mishra and G. K. Sen. Toward a theory of living state — Part I. Applied Mathematics and Computation, 56(2– 3):153–166, July 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393901204.

Meir:1994:VCF

[MS94a] A. J. Meir and Paul G.

Schmidt. A velocity-current formulation for stationary MHD flow. Applied Mathematics and Computation, 65(1–3):95–109, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901686.

Melnikov:1994:MGT

[MS94b] Yu. A. Melnikov and K. L.

Shirley. Matrices of Green’s type for the potential equation on a cylindrical surface joined to a hemisphere. Applied Mathematics and Computation, 65(1– 3):241–252, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300394901805.

Molnar:1994:VTL

[MS94c] Sándor Molnár and Ferenc Szigeti. On “verticum”-type linear systems with time-dependent linkage. Applied Mathematics and Computation, 60(1): 89–102, January 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394902089.

McArthur:1992:SGM

[MSLB92] Kelly M. McArthur, Ralph C. Smith, John Lund, and Kenneth L. Bowers. The Sinc– Galerkin method for parameterdependent self-adjoint problems. Applied Mathematics and Computation, 50(2–3): 175–202, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290125K.

Mercier:1993:AAP

[MTP93] P. Mercier, J. Teghem, and H. Pastijn. Assignment of available products to orders with the MCDM software ORESTE. Applied Mathematics and Computation, 54(2–3):183–196, March 15, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039390058M.

McKeague:1990:SCTa

[MU90a] Ian W. McKeague and Klaus J. Utikal. Stochastic calculus as a tool in survival analysis: a review. Applied Mathematics and Computation, 38(1):23–49, July 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039090079I.

McKeague:1990:SCTb

[MU90b] Ian W. McKeague and Klaus J. Utikal. Stochastic calculus as a tool in survival analysis: a review. Applied Mathematics and Computation, 39(3): 155–181, October 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090007P.

Muraskin:1991:SPD

[Mur91a] M. Muraskin. Summation-over-paths degree of freedom and the loop lattice. Applied Mathematics and Computation, 42 (2):197–207, March 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190051N.

Muratori:1991:ASP

[Mur91b] Simona Muratori. An application of the separation principle for detecting slow-fast limit cycles in a three-dimensional system. Applied Mathematics and Computation, 43(1):1–18, May 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190091Z.

Muraskin:1992:BNL

[Mur92] M. Muraskin. Beyond Newton and Leibniz. Applied Mathematics and Computation, 52(2–3): 417–438, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039290091E.

Muraskin:1993:OWO

[Mur93] M. Muraskin. Oscillations within oscillations. Applied Mathematics and Computation, 53(1): 45–82, January 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390115U.

Muraskin:1994:IOM

[Mur94] M. Muraskin. Irregular oscillations and mathematical aesthetics. Applied Mathematics and Computation, 61(2–3):183– 206, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394900477.

Muses:1994:HAH

[Mus94] C. Musès. Hypernumbers applied or how they interface with the physical world. Applied Mathematics and Computation, 60 (1):25–36, January 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394902038.

Mirchandani:1993:RNM

[MW93] Pitu B. Mirchandani and Małgorzata M. Wiecek. Routing with nonlinear multiattribute cost functions. Applied Mathematics and Computation, 54(2–3):215– 239, March 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390060R.

Nwannenna:1991:RMC

[NA91] C. C. Nwannenna and E. M.

Aghedo. Reliability monitoring and control in electric power systems. Applied Mathematics and Computation, 43 (2):117–138, May 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190028L.

Namachchivaya:1990:SBa

[Nam90a] N. Sri Namachchivaya. Stochastic bifurcation. Applied Mathematics and Computation, 38 (2):101–159, July 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300390900514.

Namachchivaya:1990:SBb

[Nam90b] N. Sri Namachchivaya. Stochastic bifurcation. Applied Mathematics and Computation, 39 (3):37–95, October 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090003L.

Nistal:1993:NIR

[NAM93] J. L. Freire Nistal, F. Alonso Amo, and F. Aguado Martín. On the naturality of the iteration and recursive specifications. Applied Mathematics and Computation, 58(1):1–18, September 15, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390042D.

Navarro:1994:CFG

[NFJ94] E. Navarro, M. V. Ferrer, and L. Jódar. Closed form general solution of nonhomogeneous implicit higher order difference systems. Applied Mathematics and Computation, 60(2–3): 113–123, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394900981.

See erratum [JNF95].

Nakamura:1994:APF

[NG94] Takashi Nakamura and Yukio-Pegio Gunji. Algebraic property of flow diagram in a timereverse cellular automaton. Applied Mathematics and Computation, 62(2–3):103–127, May 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394900779.

Newman:1992:CCI

[NP92] D. J. Newman and M. E. Primak. Complexity of circumscribed and inscribed ellipsoid methods for solving equilibrium economical models. Applied Mathematics and Computation, 52(2–3):223–231, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290079G.

Nelson:1992:IET

[NSV92] Paul Nelson, D. L. Seth, and R. Vasudevan. An integrodifferential equation for the twodimensional reflection kernel. Applied Mathematics and Computation, 49(1):1–18, May 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900534.

Niestegge:1990:AMS

[NW90] A. Niestegge and K. Witsch.

Analysis of a multigrid strokes solver. Applied Mathematics and Computation, 35(3):291– 303, February 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300390900488.

ONeil:1990:CMR

[O’N90] Kevin A. O’Neil. The continuation method for rational and deficient systems. Applied Mathematics and Computation, 37 (2):121–129, May 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090040A.

Oppenheimer:1992:ASA

[Opp92] Seth F. Oppenheimer. An abstract spectral approximation theorem from the theory of semigroups. Applied Mathematics and Computation, 47(2–3): 185–199, February 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900464.

Omari:1992:RLU

[OT92] Pierpaolo Omari and Maurizio Trombetta. Remarks on the lower and upper solutions method for second- and third-order periodic boundary value problems. Applied Mathematics and Computation, 50(1):1–21, July 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290007N.

Omari:1993:RLU

[OT93] Pierpaolo Omari and Maurizio Trombetta. Remarks on the lower and upper solutions method for second- and third-order boundary value problems. Applied Mathematics and Computation, 56(1):101, June 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390082P.

Pang:1994:DBV

[PA94] Peter Y. H. Pang and Ravi P.

Agarwal. On discrete boundary value problems arising in transport phenomena. Applied Mathematics and Computation, 60(2– 3):193–213, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901058.

Papageorgiou:1992:EWM

[Pap92] Nikolaos S. Papageorgiou. On the existence of weakly maximal programs for a multisector economy with consumption. Applied Mathematics and Computation, 48(2–3):177–186, April 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900372.

Papageorgiou:1994:SAC

[Pap94] Nikolaos S. Papageorgiou. Sensitivity analysis for a continuous time infinite horizon growth model. Applied Mathematics and Computation, 60(1): 43–53, January 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300394902054.

Patterson:1990:MRE

[Pat90] Richard L. Patterson. A Markovrenewal extension of a deterministic state-variable model of the transport and fate of a toxic chemical in the environment. Applied Mathematics and Computation, 35(3):219–230, February 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300390900433.

Patterson:1991:RAC

[Pat91] Richard L. Patterson. Regression approximations for continuoustime, stochastic compartmental models of discrete interacting populations. Applied Mathematics and Computation, 45(1): 73–89, September 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300391900896.

Patterson:1992:DUS

[Pat92] Richard L. Patterson. Distributions of unexpected services in a D-node Markov-renewal network distinguished by customer class. Applied Mathematics and Computation, 52(1):67–82, November 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039290098L.

Pattee:1993:LFM

[Pat93] H. H. Pattee. The limitations of formal models of measurement, control, and cognition. Applied Mathematics and Computation, 56(2–3):111–130, July 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390118X.

Poston:1992:BSE

[PBL92] Timothy Poston, Hyeong Ohk Bae, and Chung-Nim Lee. Bifurcation structure of the Eckalbar stock-holding model. Applied Mathematics and Computation, 48(1):21–43, March 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290016T.

Peng:1993:IPN

[Pen93] Zhang Peng. An inverse problem in nonlinear optics. Applied Mathematics and Computation, 56(1):59–70, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390079T.

Pesamosca:1991:FPA

[Pes91] Giancarlo Pesamosca. A fixed point algorithm for solving the Euclidean multifacility location problem in a tree. Applied Mathematics and Computation, 44 (2):105–120, July 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190013D.

Philos:1992:OCD

[Phi92] Ch. G. Philos. Oscillations in a class of difference equations. Applied Mathematics and Computation, 48(1):45–57, March 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290017U.

Piccardi:1992:FHO

[Pic92] Carlo Piccardi. Finite-horizon optimal control with pointwise cost functional. Applied Mathematics and Computation, 52(2– 3):345–353, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290087H.

Parent:1993:BOW

[PL93] E. Parent and F. Lebdi. Bicriterion operation of a water resources system with reliabilitybased tradeoffs. Applied Mathematics and Computation, 54(2– 3):197–213, March 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390059N.

Petkovic:1993:GOR

[PP93] Lj. Petković and M. Petković.

Generalized Ostrowski rootfinding method. Applied Mathematics and Computation, 57(2– 3):275–287, September 1, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393901525.

Philos:1994:GAN

[PPS94] Ch. G. Philos, I. K. Purnaras, and Y. G. Sficas. Global attractivity in a nonlinear difference equation. Applied Mathematics and Computation, 62 (2–3):249–258, May 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900868.

Pokhariyal:1993:AEM

[PR93] G. P. Pokhariyal and A. J. Rodrigues. An accurate epidemiological model. Applied Mathematics and Computation, 53 (1):1–12, January 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390112R.

Palusinski:1994:CEM

[PS94a] Olgierd A. Palusinski and Ferenc Szidarovszky. A Chebyshev expansion method for solving Nonlinear circuit equations. Applied Mathematics and Computation, 60(2–3):299–310, February 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901112.

Patterson:1994:NMT

[PS94b] Richard L. Patterson and David Snyder. A numerical model for tracking populations through a time-dependent stochastic network. Applied Mathematics and Computation, 61(2–3):163– 182, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900469.

Pardo:1993:APD

[PSMM93] L. Pardo, M. Salicrú, M. L. Menéndez, and D. Morales. Asymptotic properties of (r, s)- directed divergences in a stratified sampling. Applied Mathematics and Computation, 55 (2–3):131–152, May 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390018A.

Pinelis:1994:TUF

[PY94] I. Pinelis and S. Yakowitz. The time until the final zero crossing of random sums with application to nonparametric bandit theory. Applied Mathematics and Computation, 63(2–3):235–263, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039490197X.

Ping:1994:MPB

[PZ94] He Ping and Chen Zheng. A method putting bar materials to rational use and a computer program. Applied Mathematics and Computation, 63(1):59–73, June 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300394901503.

Rajendran:1992:SUT

[Raj92] K. Rajendran. Simulation of unsteady turbulent flows by quasisteady models. Applied Mathematics and Computation, 48 (1):1–11, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290014R.

Rajendran:1994:CSU

[Raj94] K. Rajendran. Computer simulation of unsteady turbulent boundary layer flow with the effect of free stream velocity. Applied Mathematics and Computation, 63(2–3):151–162, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901929.

Ramachandran:1990:ECC

[Ram90] M. P. Ramachandran. End conditions for cubic spline interpolation. Applied Mathematics and Computation, 40(2): 105–116, November 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090125M.

Raphael:1992:RTT

[Rap92] Christopher Raphael. Radiation theraphy treatment planning: an L^{2} approach. Applied Mathematics and Computation, 52(2–3): 251–277, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290082C.

Rach:1992:DES

[RBA92] R. Rach, A. Baghdasarian, and G. Adomian. Differential equations with singular coefficients. Applied Mathematics and Computation, 47(2–3): 179–184, February 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300392900453.

Reich:1994:CCV

[RBR94] Robin M. Reich, Charles D. Bonham, and Kimberly K. Remington. Comparison of classical vs. inverse sampling for estimating plant cover using Monte Carlo simulations. Applied Mathematics and Computation, 64(2–3): 171–190, September 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900620.

Ranasinghe:1991:SBE

[RC91] A. I. Ranasinghe and M. H.

Chang. Solution of the Burgers equation on semiinfinite and finite intervals via a stream function. Applied Mathematics and Computation, 41(2): 145–158, January 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190067W.

Reddy:1990:NIM

[Red90a] Y. N. Reddy. A numerical integration method for solving singular perturbation problems. Applied Mathematics and Computation, 37(2):83–95, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300390900374.

Reddy:1990:NTC

[Red90b] Y. N. Reddy. Numerical techniques for a class of singular perturbation problems. Applied Mathematics and Computation, 35(1):1–11, January 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090089L.

Reddy:1991:NNI

[Red91] Y. N. Reddy. A note on the numerical-integration method for solving singular perturbation problems. Applied Mathematics and Computation, 43 (2):175–180, May 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190032I.

Regan:1994:LHW

[Reg94]  Marian P. Regan.  Linear

half-wave rectification of modulated sinusoids. Applied Mathematics and Computation, 62 (1):61–79, April 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901333.

Reuster:1993:TDC

[RK93] D. Reuster and M. Kaye. A technique for developing complete orthonormal basis sets using general solutions of Bessel’s differential equation. Applied Mathematics and Computation, 55 (2–3):255–264, May 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300393900249.

Rosen:1993:MM

[Ros93]  Robert Rosen.  On models

and modeling. Applied Mathematics and Computation, 56 (2–3):359–372, July 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393901282.

Rossler:1993:EA

[RR93] Otto E. Rossler and Jonas O.

Rossler. The endo approach. Applied Mathematics and Computation, 56(2–3):281–287, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390125X.

Razzaghi:1990:SCI

[RRA90] Mohsen Razzaghi, Mehdi Razzaghi, and Abdollah Arabshahi. Solutions of convolution integral and Fredholm integral equations via double Fourier series. Applied Mathematics and Computation, 40(3):215–224, December 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0096300308800033.

Rassias:1992:OPC

[RS92] Themistocles M. Rassias and H. M. Srivastava. The orthogonality property of the classical Laguerre polynomials. Applied Mathematics and Computation, 50(2–3):167–173, March 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290124J.

Rehbock:1992:EPF

[RTJL92] V. Rehbock, K. L. Teo, L. S. Jennings, and C. S. Lee. An exact penalty function approach to alltime-step constrained discretetime optimal control problems. Applied Mathematics and Computation, 49(2–3):215–230, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039290025V.

Ruan:1992:ASV

[Rua92] Shigui Ruan. Asymptotic stability for Volterra intergrodifferential systems. Applied Mathematics and Computation, 52(2– 3):207–222, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290078F.

Ruggiero:1993:PPV

[Rug93] Valeria Ruggiero. Polynomial preconditioning on vector computers. Applied Mathematics and Computation, 59(2–3):131– 150, December 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/009630039390085S.

Russell:1993:FVD

[RW93] David L. Russell and Luther W.

White. Formulation and validation of dynamical models for narrow plate motion. Applied Mathematics and Computation, 58(2– 3):103–141, October 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390133Y.

Rashid:1994:TLF

[RWB94] Abdur Rashid, Cao Weiming, and Guo Benyu. Three level Fourier spectral approximations for fluid flow with low Mach number. Applied Mathematics and Computation, 63(2– 3):131–149, July 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901910.

Sheng:1994:NVP

[SA94] Qin Sheng and Ravi P. Agarwal. Nonlinear variation of parameter methods for summary difference equations in several independent variables. Applied Mathematics and Computation, 61(1):39–60, March 1, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901457.

Sage:1991:IIC

[Sag91] Andrew P. Sage. Invariant imbedding in control, estimation, and system identification. Applied Mathematics and Computation, 45(2):99–119, September 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190074W.

Sarper:1993:EAN

[Sar93] Hüseyin Sarper. Evaluation of the accuracy of Naslund’s approximation for managerial decision making under uncertainty. Applied Mathematics and Computation, 55(1):73–87, April 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390006Z.

Savageau:1993:FMR

[Sav93] Michael A. Savageau. Finding multiple roots of nonlinear algebraic equations using S-system methodology. Applied Mathematics and Computation, 55(2– 3):187–199, May 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390020F.

Suleiman:1992:UNI

[SB92] Mohamed Suleiman and Saiman Baok. Using nonconvergence of iteration to partition ODEs. Applied Mathematics and Computation, 49(2–3):111–140, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900202.

Schulz:1991:SSM

[Sch91a] Lech Schulz. Set-spectrum method in structure comparison and the concept of topological kinship. Applied Mathematics and Computation, 41 (1):1–20, January 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190103T.

Schumitzky:1991:NEA

[Sch91b] Alan Schumitzky. Nonparametric EM Algorithms for estimating prior distributions. Applied Mathematics and Computation, 45(2):143–157, September 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190077Z.

Schmid:1992:AAC

[Sch92] Ulrich Schmid. On the asymptotics of the average CRI length of the slotted ALOHA collision resolution algorithm. Applied Mathematics and Computation, 48(1):59–70, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039290018V.

Selvakumar:1994:CMS

[Sel94] K. Selvakumar. A computational method for solving singular perturbation problems using exponentially fitted finite difference schemes. Applied Mathematics and Computation, 66(2– 3):277–292, December 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901236.

Semper:1992:NCO

[Sem92] Bill Semper. A note on the computation of the outflow derivative for singular perturbation problems. Applied Mathematics and Computation, 51(2–3): 135–141, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290069D.

Sengupta:1990:JIS

[Sen90] A. Sengupta. Jordan invariant subspaces via algebraic generalized inverse. Applied Mathematics and Computation, 38 (2):75–85, July 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300390900499.

Sengupta:1993:NRC

[Sen93] A. Sengupta. A note on a reduction of Cauchy singular integral equation to Fredholm equation in L_{p}. Applied Mathematics and Computation, 56(1):97–100, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039390081O.

Sen:1992:DFP

[SH92] S. K. Sen and Gary W. Howell.

Direct fail-proof triangularization algorithms for AX + XB = C with Error-Free and parallel implementations. Applied Mathematics and Computation, 50(2– 3):255–278, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290128N.

Sharp:1990:SDEa

[Sha90a] Keith P. Sharp. Stochastic differential equations in finance. Applied Mathematics and Computation, 37(2):131–148, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090041Z.

Sharp:1990:SDEb

[Sha90b] Keith P. Sharp. Stochastic differential equations in finance. Applied Mathematics and Computation, 39(3):207–224, October 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039090009R.

Shalaby:1994:MER

[Sha94] M. A. Shalaby. Minimal eigenvalue of a real symmetric positive definite Toeplitz matrix. Applied Mathematics and Computation, 64(2–3):93–99, September 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394900558.

Shetty:1990:ASL

[She90] Bala Shetty. Approximate solutions to large scale capacitated facility location problems. Applied Mathematics and Computation, 39(2):159–175, September 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300390900293.

Shimizu:1993:BAS

[Shi93] H. Shimizu. Biological autonomy: the self-creation of constraints. Applied Mathematics and Computation, 56(2–3):177–201, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039390122U.

Stokes:1994:PMM

[SHR94] M. L. Stokes, D. H. Huddleston, and M. G. Remotigue. A practical model for multidisciplinary analysis data and algorithm abstraction. Applied Mathematics and Computation, 65(1–3):111– 123, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901694.

Simos:1992:EAS

[Sim92] T. E. Simos. An explicit almost P -stable two-step method with phase-lag of order infinity for the numerical integration of second-order Pacific initialvalue problems. Applied Mathematics and Computation, 49(2– 3):261–268, June 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290029Z.

Simos:1994:SNV

[Sim94] T. E. Simos. Some new variablestep methods with minimal phase lag for the numerical integration of special second-order initialvalue problem. Applied Mathematics and Computation, 64 (1):65–72, August 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901392.

Sivaloganathan:1992:FEP

[Siv92] S. Sivaloganathan. A finiteelement, parameter stepping solution procedure for the steadystate Boussinesq equations. Applied Mathematics and Computation, 51(1):43–83, September 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039290134M.

Shang:1994:NSD

[SKCW94] H. M. Shang, Y. M. Kim, C. P. Chen, and T. S. Wang. Numerical studies of droplet-turbulence interactions. Applied Mathematics and Computation, 65(1–3): 63–78, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039490166X.

Slater:1991:PBM

[Sla91] Paul B. Slater. Physical-based models of internal migration. Applied Mathematics and Computation, 43(1):95–103, May 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300391900966.

Slater:1993:IMA

[Sla93a] Paul B. Slater. International migration and air travel: global smoothing and estimation. Applied Mathematics and Computation, 53(2–3):225–234, February 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039390103L.

Slater:1993:WPD

[Sla93b] Paul B. Slater. World population distribution: smoothed representations. Applied Mathematics and Computation, 53(2–3): 207–223, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390102K.

Szidarovszky:1994:LDP

[SM94] Ferenc Szidarovszky and Sàndor Molnàr. Learning in a dynamic producer–consumer market. Applied Mathematics and Computation, 62(2–3):223–233, May 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900841.

Szidarovszky:1991:PSL

[SMO91] Ferenc Szidarovszky, Sándor Molnár, and Koji Okuguchi. An N-person Stackelberg leader– leader model. Applied Mathematics and Computation, 46(3): 221–232, December 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391901187.

Seda:1992:PBV

[ŠNG92] V. Šeda, Juan J. Nieto, and M. Gera. Periodic boundary value problems for nonlinear higher order ordinary differential equations. Applied Mathematics and Computation, 48 (1):71–82, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290019W.

Szidarovszky:1990:DOM

[SO90] Ferenc Szidarovszky and Koji Okuguchi. Dynamic oligopoly: models with incomplete information. Applied Mathematics and Computation, 38(2): 161–177, July 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300390900525.

Soni:1993:EGG

[Son93] Bharat K. Soni. Elliptic grid generation system: control functions revisited — I. Applied Mathematics and Computation, 59(2– 3):151–163, December 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390086T.

Soriano:1993:GMP

[Sor93] J. M. Soriano. Global minimum point of a convex function. Applied Mathematics and Computation, 55(2–3):213–218, May 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393900227.

Soriano:1994:EPC

[Sor94] J. M. Soriano. Extremum points of a convex function. Applied Mathematics and Computation, 66(2–3):261–266, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039490121X.

See erratum [Sor97].

Soriano:1997:EEP

[Sor97] J. M. Soriano. Erratum: “Extremum points of a convex function”. Applied Mathematics and Computation, 84(1):95, June 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0096300397002105.

See [Sor94].

Sosnowska:1991:CKW

[Sos91] Honorata Sosnowska. Controllability in the Kornai–Weibull queueing model of a market in a shortage economy. Applied Mathematics and Computation, 43(3):207–220, June 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300391900069.

Szpiro:1993:PSC

[SP93] A. A. Szpiro and V. Protopopescu. Parabolic systems with competitive interactions and control on initial conditions. Applied Mathematics and Computation, 59(2–3):215–245, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300393900902.

Szidarovszky:1994:SME

[SP94] Ferenc Szidarovszky and Olgierd A. Palusinski. A special matrix equation and its application in Microelectronics. Applied Mathematics and Computation, 64(2–3):115–119, September 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300394900574.

Soyeux:1993:BQA

[SPD93] A. Soyeux, E. Parent, and L. Duckstein. A biattribute quality analysis of the food industry. Applied Mathematics and Computation, 58(2–3): 199–214, October 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300393901363.

Spitaleri:1990:MMG

[Spi90] R. M. Spitaleri. A multigrid method for grid generation with line-spacing control. Applied Mathematics and Computation, 40(2):125–134, November 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090127O.

Sengupta:1990:ASL

[SR90] A. Sengupta and M. Rama Mohana Rao. On approximate solution of linear singular integral equations. Applied Mathematics and Computation, 36 (3):147–161, April 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090010Z.

Srivastava:1994:ORG

[Sri94] H. M. Srivastava. Orthogonality relations and generating functions for the generalized Bessel polynomials. Applied Mathematics and Computation, 61(2–3): 99–134, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900426.

Sanchez:1991:DFD

[SS91] Paul J. Sanchez and Susan M.

Sanchez. Design of frequencydomain experiments for discretevalued factors. Applied Mathematics and Computation, 42 (1):1–21, March 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190111Y.

Sivasundaram:1992:AIA

[SS92] S. Sivasundaram and Yong Sun. Application of interval analysis to impulsive differential equations. Applied Mathematics and Computation, 47(2–3): 201–210, February 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900475.

Shahruz:1994:ASL

[SS94a] S. M. Shahruz and A. L. Schwartz. An approximate solution for linear boundary-value problems with slowly varying coefficients. Applied Mathematics and Computation, 60(2–3): 285–298, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901104.

Shivaji:1994:SID

[SS94b] Ratnasingham Shivaji and Bharat Soni. Special issue on differential equations and computational simulations I. Applied Mathematics and Computation, 65(1– 3):1–2, September/October 1994.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901619.

Steit:1994:CSQ

[Ste94] H. A. H. Abou Steit. Cubic spline on quintic spline method for nonlinear boundary value problems. Applied Mathematics and Computation, 66(2–3): 267–275, December 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901228.

Serrano:1990:REEa

[SU90a] Sergio E. Serrano and T. E.

Unny. Random evolution equations in hydrology. Applied Mathematics and Computation, 38 (3):201–226, August 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090023V.

Serrano:1990:REEb

[SU90b] Sergio E. Serrano and T. E.

Unny. Random evolution equations in hydrology. Applied Mathematics and Computation, 39 (3):97–122, October 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090004M.

Srinivasu:1992:QCN

[SV92a] P. D. N. Srinivasu and M. Venkatesulu. Quadratically convergent numerical schemes for nonstandard initial value problems. Applied Mathematics and Computation, 47(2–3):145–154, February 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039290043Z.

Srinivasu:1992:SNT

[SV92b] P. D. N. Srinivasu and M. Venkatesulu. Solutions of nonstandard n th order initial value problems. Applied Mathematics and Computation, 47(2–3): 155–178, February 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300392900442.

Saxena:1994:STP

[SV94] Anjula Saxena and Ezio Venturino. Solving two-point boundary value problems by means of deficient quartic splines. Applied Mathematics and Computation, 66(1):25–40, November 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300394901562.

Silliman:1991:GRF

[SW91] S. E. Silliman and A. L. Wright. Generation of random fields characterized by discrete regions of constant value. Applied Mathematics and Computation, 45(3): 293–311, October 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391900852.

Shaw:1992:ATG

[SW92] J. A. Shaw and N. P. Weatherill. Automatic topology generation for multiblock grids. Applied Mathematics and Computation, 52(2–3):355–388, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290088I.

Sheng:1994:IAN

[SW94] Zhang Lian Sheng and M. A.

Wolfe. An interval algorithm for nondifferentiable global optimization. Applied Mathematics and Computation, 63(2– 3):101–122, July 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901899.

Sworder:1991:HAC

[Swo91] D. D. Sworder. Hybrid adaptive control. Applied Mathematics and Computation, 45(2): 173–192, September 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300391900793.

Szidarovszky:1993:CDD

[SY93] Ferenc Szidarovszky and Jerome Yen. On the controllability of discrete dynamic oligopolies under adaptive expectations. Applied Mathematics and Computation, 56(1):49–57, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390078S.

Shih:1994:IGG

[SYS94] Ming H. Shih, Tzu-Yi Yu, and Bharat K. Soni. Interactive grid generation and NURBS applications. Applied Mathematics and Computation, 65(1–3):279–294, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039490183X.

Szidarovszky:1990:CNA

[Szi90] Ferenc Szidarovszky. On the convergence of nonstationary algorithms modeled by point-to-set maps. Applied Mathematics and Computation, 39(2):107– 110, September 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090025X.

Tabatabai:1993:RET

[TA93] M. A. Tabatabai and I. K. Argyros. Robust estimation and testing for general nonlinear regression models. Applied Mathematics and Computation, 57 (1):85–101, August 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393900146.

Tsai:1992:CAM

[TCS92] J. S. H. Tsai, C. M. Chen, and L. S. Shieh. A computeraided method for solvents and spectral factors of matrix polynomials. Applied Mathematics and Computation, 47(2–3): 211–235, February 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900486.

Tecle:1994:IMP

[TDK94] Aregai Tecle, Lucien Duckstein, and Pekka Korhonen. Interactive, multiobjective programming for forest resources management. Applied Mathematics and Computation, 63(1):75–93, June 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901511.

Tennakone:1990:SBC

[Ten90] K. Tennakone. Stability of the biomass-carbon dioxide equilibrium in the atmosphere: mathematical model. Applied Mathematics and Computation, 35(2): 125–130, January 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090113H.

Tennakone:1991:KGA

[Ten91] K. Tennakone. The kinetics of the growth of an accidentally created chiral biomolecule in the racemic prebiotic medium. Applied Mathematics and Computation, 43(2):139–144, May 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190029M.

Tesfatsion:1991:I

[Tes91a] Leigh Tesfatsion. Introduction. Applied Mathematics and Computation, 45(2):89–90, September 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039190071T.

Tesfatsion:1991:WRKa

[Tes91b] Leigh Tesfatsion. Work by Robert Kalaba on automated sensitivity analysis. Applied Mathematics and Computation, 45(2):121–133, September 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/009630039190075X.

Tesfatsion:1991:WRKb

[Tes91c]  Leigh Tesfatsion.  Work by

Robert Kalaba on multicriteria estimation. Applied Mathematics and Computation, 45(2): 159–171, September 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391900782.

Teo:1993:SFS

[TFM93] K. L. Teo, M. E. Fisher, and J. B.

Moore. A suboptimal feedback stabilizing controller for a class of nonlinear regulator problems. Applied Mathematics and Computation, 59(1):1–17, November 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S009630039390038G.

Tapia:1991:CML

[TGK91] M. A. Tapia, T. A. Guima, and A. Katbab. Calculus for a multivalued-logic algebraic system. Applied Mathematics and Computation, 42 (3):255–285, April 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391900047.

Todorov:1994:EHG

[TH94] Pavel G. Todorov and John H. Hubbell. On the Epstein– Hubbell generalized elliptic-type integral. Applied Mathematics and Computation, 61(2–3): 157–161, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394900450.

Teo:1990:NCD

[TLG90] K. L. Teo, Y. Liu, and C. J. Goh. Nonlinearly constrained discretetime optimal-control problems. Applied Mathematics and Computation, 38(3):227–248, August 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039090024W.

Trigiante:1992:BLS

[TM92] D. Trigiante and K. N. Murty. On best least square solutions of two-point boundary value problems via modified Gram–Schmidt process. Applied Mathematics and Computation, 51(2–3): 105–118, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290067B.

Thompson:1990:ONEa

[TT90a] M. E. Thompson and A. Thavaneswaran. Optimal nonparametric estimation for some semimartingale stochastic differential equations. Applied Mathematics and Computation, 37 (3):169–183, June 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300390900547.

Thompson:1990:ONEb

[TT90b] M. E. Thompson and A. Thavaneswaran. Optimal nonparametric estimation for some semimartingale stochastic differential equations. Applied Mathematics and Computation, 39(3): 123–137, October 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090005N.

Tennakone:1990:BCS

[TW90] K. Tennakone and S. Wickramanayaka. Biochemical chiral selection in the presence of instabilities, chaos, and noise. Applied Mathematics and Computation, 38(1):1–6, July 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090077G.

Udwadia:1991:NPC

[Udw91] Firdaus E. Udwadia. Noncollocated point control of nondispersive distributed-parameter systems using time delays. Applied Mathematics and Computation, 42(1):23–63, March 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190112Z.

Udwadia:1992:NCD

[Udw92a] Firdaus E. Udwadia. Noncollocated control of distributedparameter nondispersive systems with tip inertias using time delays. Applied Mathematics and Computation, 47 (1):47–75, January 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300392900648.

Udwadia:1992:SCR

[Udw92b] Firdaus E. Udwadia. Some convergence results related to the JOR iterative method for symmetric, positive-definite matrices. Applied Mathematics and Computation, 47(1):37–45, January 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300392900637.

Ueno:1991:SAC

[Uen91] Sueo Ueno. Stochastic approach to the Cauchy problem of linearized Couette flow. Applied Mathematics and Computation, 45(2):207–217, September 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190081W.

Udwadia:1994:CIM

[UK94] Firdaus E. Udwadia and Ravi Kumar. Convergence of iterative methods for nonclassically damped dynamic systems. Applied Mathematics and Computation, 61(1):61–97, March 1, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901465.

Unny:1990:ISI

[Unn90] T. E. Unny. Introduction to special issue on theory and application of stochastic differential equations. Applied Mathematics and Computation, 39 (3):1–2, October 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090001J.

Vassilevski:1991:MPE

[Vas91] P. S. Vassilevski. Multilevel preconditioners for elliptic problems by substructuring. Applied Mathematics and Computation, 46(2): 79–106, November 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190018I.

Voulov:1992:ASH

[VB92] H. D. Voulov and D. D. Bainov. Asymptotic stability for a homogeneous singularly perturbed system of differential equations with unbounded delay. Applied Mathematics and Computation, 51(1):1–16, September 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290132K.

Verschelde:1991:NSS

[VBH91] Jan Verschelde, Marc Beckers, and Ann Haegemans. A new start system for solving deficient polynomial systems using continuation. Applied Mathematics and Computation, 44(3): 225–239, August 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190059V.

Varadarajan:1991:RCT

[VCM91] V. Varadarajan, W. H. Choe, and G. H. Miley. Rapid calculation of tokamak equilibria with toroidal rotation via minimization of a functional. Applied Mathematics and Computation, 43(2):145–163, May 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190030Q.

Veverka:1992:MRM

[Vev92] Vladimír Veverka. A method of reconciliation of measured data with nonlinear constraints. Applied Mathematics and Computation, 49(2–3):141–176, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290021R.

Voller:1992:MEN

[Vol92] Rudolf L. Voller. Monotone enclosure for nonlinear PDEs using multigrid techniques. Applied Mathematics and Computation, 47(2–3):121–135, February 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290041X.

Vazirinejad:1990:CSE

[VS90] Shamsedin Vazirinejad and Soroosh Sorooshian. A Comparison of simultaneous-equations, weighted regression, and noise-in-variables models. Applied Mathematics and Computation, 40(3): 187–202, December 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S009630030880001X.

Vulanovic:1990:SIN

[Vul90]  Relja Vulanović.  Some im-

provements of the nonequidistant Engquist–Osher scheme. Applied Mathematics and Computation, 40(2):147–164, November 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039090129Q.

Vu:1994:NSF

[VWSS94] Bruce Vu, Ten-See Wang, Ming-Hsin Shih, and Bharat Soni. Navier–Stokes flow field analysis of compressible flow in a high pressure safety relief valve. Applied Mathematics and Computation, 65(1–3):345–353, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901872.

Waissi:1990:NSC

[Wai90] Gary R. Waissi. A new O(n^{2}) shortest chain algorithm. Applied Mathematics and Computation, 37(2):111–120, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300390900396.

Wampler:1992:BNC

[Wam92] Charles W. Wampler. Bezout number calculations for multi-homogeneous polynomial systems. Applied Mathematics and Computation, 51(2–3): 143–157, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290070H.

Wang:1992:SIT

[Wan92] Shao-Ming Wang. Spline interpolations over type-2 triangulations. Applied Mathematics and Computation, 49(2–3): 299–313, June 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290031U.

Wang:1993:RNN

[Wan93] Jun Wang. A recurrent neural network for real-time matrix inversion. Applied Mathematics and Computation, 55 (1):89–100, April 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393900072.

Wang:1994:CSA

[Wan94] Shao-Ming Wang. A cubic structure on an arbitrary quadrilateral. Applied Mathematics and Computation, 60(1): 55–68, January 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394902062.

Warren:1991:TTS

[War91] Richard H. Warren. Tours for the traveling salesman. Applied Mathematics and Computation, 46(3):167–177, December 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391901154.

Warren:1994:SCT

[War94] Richard H. Warren. Special cases of the traveling salesman problem. Applied Mathematics and Computation, 60(2–3): 171–177, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901031.

Wazwaz:1993:NSG

[Waz93] Abdul-Majid Wazwaz. On the numerical solution of the Goursat problem. Applied Mathematics and Computation, 59(1): 89–95, November 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390036E.

Werbos:1993:CSF

[Wer93] Paul J. Werbos. Chaotic solitons and the foundations of physics: a potential revolution. Applied Mathematics and Computation, 56(2–3):289–339, July 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390126Y.

White:1991:EEP

[Whi91] L. W. White. Estimation of elastic parameters in a nonlinear elliptic model of a plate. Applied Mathematics and Computation, 42(2):139–187, March 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190049S.

White:1993:EMP

[Whi93] L. W. White. Estimation of material parameters in a dynamic nonlinear plate model. Applied Mathematics and Computation, 54(1):25–63, March 1, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390032A.

Wong:1992:SAS

[WJ92] Yau Shu Wong and Xingzhi Ji. On shooting algorithm for Sturm–Liouville eigenvalue problems with periodic and semi-periodic boundary conditions. Applied Mathematics and Computation, 51(2–3): 87–104, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290066A.

Wexler:1991:NMT

[WM91] Anthony S. Wexler and Donald J. Marsh. Numerical methods for three-dimensional models of the urine concentrating mechanism. Applied Mathematics and Computation, 45(2): 219–240, September 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190082X.

Weatherill:1991:SPF

[WN91] N. P. Weatherill and D. K. Natakusumah. The simulation of potential flow around multiple bodies using overlapping connected meshes. Applied Mathematics and Computation, 46 (1):1–21, November 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391900988.

Wu:1994:PRC

[Wu94] Berlin Wu. Pattern recognition and classification in time series analysis. Applied Mathematics and Computation, 62 (1):29–45, April 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901317.

Xiguan:1994:FPS

[XYY94] Lu Xiguan, Li Yong, and Su Yi. Finding periodic solutions of ordinary differential equations via homotopy method. Applied Mathematics and Computation, 62(2–3):153–170, May 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394900795.

Yakowitz:1992:EPA

[Yak92a] Diana Yakowitz. An exact penalty algorithm for recourseconstrained stochastic linear programs. Applied Mathematics and Computation, 49(1):39–62, May 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300392900556.

Yakowitz:1992:DMM

[Yak92b] Sid Yakowitz. A decision model and methodology for the AIDS epidemic. Applied Mathematics and Computation, 52(1):149– 172, November 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392901038.

Yang:1994:PAA

[Yan94] Hailang Yang. A projective algorithm for approximation of fractal sets. Applied Mathematics and Computation, 63(2– 3):201–212, July 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300394901953.

Yee:1993:ADM

[Yee93] Eugene Yee. Application of the decomposition method to the solution of the reaction-convection-diffusion equation. Applied Mathematics and Computation, 56(1):1–27, June 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039390075P.

Yang:1991:MNA

[YF91] Jun Yang and Changxin Fan. Modeling of nonlinear and active cochlear mechanics and its Volterra-series analysis. Applied Mathematics and Computation, 43(3):241–263, June 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190010K.

Yakowitz:1990:CAM

[YGH90] S. Yakowitz, J. Gani, and R. Hayes. Cellular automaton modeling of epidemics. Applied Mathematics and Computation, 40(1):41–54, November 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039090097M.

Yip:1991:IFD

[Yip91] Paul Yip. Inference on a fatal disease. Applied Mathematics and Computation, 41(1):

89–97, January 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190109Z.

Yip:1992:IPS

[Yip92] Paul Yip. An inference procedure for a simple epidemic model: a martingale approach. Applied Mathematics and Computation, 47(1):77–89, January 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900659.

Yip:1993:SIP

[Yip93] Paul Yip. Statistical inference procedure for a hypergeometric model for capture–recapture experiments. Applied Mathematics and Computation, 58(2–3): 225–240, October 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300393901385.

Yakowitz:1993:MAD

[YLS93] D. S. Yakowitz, L. J. Lane, and F. Szidarovszky. Multi-attribute decision making: dominance with respect to an importance order of the attributes. Applied Mathematics and Computation, 54(2– 3):167–181, March 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/009630039390057L.

Yang:1994:SAG

[YS94] Jiann-Cherng Yang and Bharat K. Soni. Structured adaptive grid generation. Applied Mathematics and Computation, 65(1–3):265– 278, September/October 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901821.

Zaghrout:1991:OSO

[ZA91] A. A. S. Zaghrout and Z. Ali. On oscillations of second order linear differential equations. Applied Mathematics and Computation, 46(2):135–144, November 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039190021E.

Zaghrout:1991:SPF

[Zag91] Afaf A. S. Zaghrout. Stability and persistence of facultative mutualism with populations interacting in a food chain. Applied Mathematics and Computation, 45(1):1–15, September 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300391900863.

Zaghrout:1992:ABS

[Zag92] Afaf A. S. Zaghrout. Asymptotic behavior of solutions of competition in gradostat with two limiting complementary substrates. Applied Mathematics and Computation, 49(1):19–37, May 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300392900545.

Zaslavsky:1993:AAM

[Zas93] Leonid Yu. Zaslavsky. An adaptive algebraic multigrid for multigroup neutron diffusion reactor core calculations. Applied Mathematics and Computation, 53 (1):13–26, January 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039390113S.

Zaghrout:1991:GGB

[ZED91] A. A. S. Zaghrout and N. El-Dawaliby. Generalizations of Gronwall–Bellman–Reid inequalities. Applied Mathematics and Computation, 41(2):129– 143, January 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039190066V.

Zufiria:1990:CMF

[ZG90] Pedro J. Zufiria and Ramesh S.

Guttalu. A computational method for finding all the roots of a vector function. Applied Mathematics and Computation, 35(1):13–59, January 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090090P.

Zheng:1993:MMC

[Zhe93] Chen Zheng. A mathematical model and computer program of economic evaluation and prediction. Applied Mathematics and Computation, 59(1): 63–72, November 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S009630039390041C.

Zhu:1994:SMF

[Zhu94] Jianping Zhu. Simulating multiphase flow in porous media on KSR-1 parallel computers. Applied Mathematics and Computation, 65(1–3):19–31, September/October 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901635.

Zouyousefain:1990:SRD

[ZL90] M. Zouyousefain and S. Leela. Stability results for difference equations of Volterra type. Applied Mathematics and Computation, 36(1):51–61, March 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090074D.

Zhang:1991:GMV

[ZM91] Hong Zhang and Xiao-Yi Min.

On the Galerkin method with vector basis functions. Applied Mathematics and Computation, 42(1):85–103, March 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300391901143.

Zheng:1992:UCS

[ZP92] Chen Zheng and He Ping. Using computers to seek the optimum combination of investment. Applied Mathematics and Computation, 49(2–3):197–203, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630039290023T.

Zuhe:1990:BNP

[ZW90a] Shen Zuhe and M. A. Wolfe. A ball Newton point algorithm for bounding zeros of analytic functions. Applied Mathematics and Computation, 36(1):1–14, March 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630039090071A.

Zuhe:1990:IEU

[ZW90b] Shen Zuhe and M. A. Wolfe. On interval enclosures using slope arithmetic. Applied Mathematics and Computation, 39(1):89– 105, September 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039090124L.

Zuhe:1992:STN

[ZW92] Shen Zuhe and M. A. Wolfe. Slope tests for Newton-type methods. Applied Mathematics and Computation, 52(2–3): 403–416, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630039290090N.

Zhihong:1994:VPF

[ZXY94] Guan Zhihong, Wen Xiangcai, and Liu Yongqing. Variation of the parameters formula and the problem of BIBO for singular measure differential systems with impulse effect. Applied Mathematics and Computation, 60(2– 3):153–169, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300394901023.